### v4

# A contramodule generalization of Neeman’s flat and projective module theorem

Abstract.
This paper builds on top of [34]. We consider a complete, separated topological ring $\mathfrak{R}$ with a countable base of neighborhoods of zero consisting of open two-sided ideals. The main result is that the homotopy category of projective left $\mathfrak{R}$-contramodules is equivalent to the derived category of the exact category of flat left $\mathfrak{R}$-contramodules, and also to the homotopy category of flat cotorsion left $\mathfrak{R}$-contramodules. In other words, a complex of flat $\mathfrak{R}$-contramodules is contraacyclic (in the sense of Becker) if and only if it is an acyclic complex with flat $\mathfrak{R}$-contramodules of cocycles, and if and only if it is coacyclic as a complex in the exact category of flat $\mathfrak{R}$-contramodules. These are contramodule generalizations of theorems of Neeman and of Bazzoni, Cortés-Izurdiaga, and Estrada.
Introduction
The flat and projective periodicity theorem of Benson and Goodearl [6, Theorem 2.5] claims that, in the context of modules over an arbitrary ring $R$, if $0\longrightarrow F\longrightarrow P\longrightarrow F\longrightarrow 0$ is a short exact sequence with a flat module $F$ and a projective module $P$, then the module $F$ is actually projective. This result was rediscovered by Neeman [26, Remark 2.15], who proved the following much stronger claims.
Let $F^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be a complex of flat left $R$-modules. Then the following three conditions are equivalent:
  1. (i)
    for every complex of projective left $R$-modules $P^{\text{\smaller\smaller$\scriptstyle\bullet$}}$, any morphism of complexes of $R$-modules $P^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow F^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero;
  2. (ii)
    $F^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is an acyclic complex of $R$-modules with flat $R$-modules of cocycles;
  3. (iii)
    $F^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is a directed colimit of contractible complexes of projective left $R$-modules.
See [26, Theorem 8.6].
A further development came in the paper by Bazzoni, Cortés-Izurdiaga, and Estrada [2], where the cotorsion periodicity theorem [2, Theorem 1.2(2), Proposition 4.8(2), or Theorem 5.1(2)] was proved. It claims that if $0\longrightarrow M\longrightarrow C\longrightarrow M\longrightarrow 0$ is a short exact sequence of $R$-modules and the $R$-module $C$ is cotorsion, then the $R$-module $M$ is cotorsion, too. As a corollary, it was essentially shown in [2, Theorem 5.3] that the following two conditions on a complex of flat left $R$-modules $F^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ are equivalent to (i–iii):
  1. (iv)
    for every complex of cotorsion left $R$-modules $C^{\text{\smaller\smaller$\scriptstyle\bullet$}}$, any morphism of complexes of $R$-modules $F^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow C^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero;
  2. (v)
    for every complex of flat cotorsion left $R$-modules $G^{\text{\smaller\smaller$\scriptstyle\bullet$}}$, any morphism of complexes of $R$-modules $F^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow G^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero.
The aim of this paper is to prove the following generalization of the theorems of Neeman and of Bazzoni, Cortés-Izurdiaga, and Estrada. Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open two-sided ideals. Let $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be a complex of flat left $\mathfrak{R}$-contramodules. Then the following six conditions are equivalent:
  1. (ic)
    for every complex of projective left $\mathfrak{R}$-contamodules $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$, any morphism of complexes of $\mathfrak{R}$-contramodules $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero;
  2. (iic)
    $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is an acyclic complex of $\mathfrak{R}$-contramodules with flat $\mathfrak{R}$-contramodules of cocycles;
  3. (iiic)
    $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ can be obtained from contractible complexes of flat $\mathfrak{R}$-contramodules using extensions and directed colimits;
  4. (ivc)
    $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is an $\aleph_{1}$-directed colimit of total complexes of short exact sequences of complexes of countably presentable flat $\mathfrak{R}$-contramodules;
  5. (vc)
    for every complex of cotorsion left $\mathfrak{R}$-contramodules $\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$, any morphism of complexes of $\mathfrak{R}$-contramodules $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero;
  6. (vic)
    for every complex of flat cotorsion left $\mathfrak{R}$-contramodules $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$, any morphism of complexes of $\mathfrak{R}$-contramodules $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero.
The equivalence of the three conditions (iic) $\Longleftrightarrow$ (iiic) $\Longleftrightarrow$ (ivc) was established in the paper [34, Theorem 13.2]. The aim of the present paper is to prove the equivalence of the four conditions (ic) $\Longleftrightarrow$ (iic) $\Longleftrightarrow$ (vc) $\Longleftrightarrow$ (vic).
The study of arbitrary (unbounded) complexes of projective modules goes back to Jørgensen’s paper [19]. The terminology “contraderived category” was introduced by the present author in [27, 28], inspired by Keller’s terminology “coderived category” in [21] (see also [22]). The contemporary point of view on the topic came with Becker’s paper [4, Proposition 1.3.6]. Let us spell out some definitions.
Let $\mathsf{B}$ be an abelian (or exact) category with enough projective objects. A complex $B^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ in $\mathsf{B}$ is said to be contraacyclic (in the sense of Becker) if, for every complex of projective objects $P^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ in $\mathsf{B}$, any morphism of complexes $P^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow B^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero. The triangulated Verdier quotient category $\mathsf{D}^{\mathsf{bctr}}(\mathsf{B})=\mathsf{Hot}(\mathsf{B})/\mathsf{Ac}^{\mathsf{bctr}}(\mathsf{B})$ of the homotopy category $\mathsf{Hot}(\mathsf{B})$ of complexes in $\mathsf{B}$ by the thick subcategory of contraaacyclic complexes $\mathsf{Ac}^{\mathsf{bctr}}(\mathsf{B})\subset\mathsf{Hot}(\mathsf{B})$ is called the (Becker) contraderived category of the abelian/exact category $\mathsf{B}$.
Dually, let $\mathsf{A}$ be an exact category with enough injective objects. A complex $A^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ in $\mathsf{A}$ is said to be coacyclic (in the sense of Becker) if, for every complex of injective objects $J^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ in $\mathsf{A}$, any morphism of complexes $A^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow J^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero. The quotient category $\mathsf{D}^{\mathsf{bco}}(\mathsf{A})=\mathsf{Hot}(\mathsf{A})/\mathsf{Ac}^{\mathsf{bco}}(\mathsf{A})$ of the homotopy category $\mathsf{Hot}(\mathsf{A})$ by the thick subcategory of coacyclic complexes $\mathsf{Ac}^{\mathsf{bco}}(\mathsf{A})\subset\mathsf{Hot}(\mathsf{A})$ is called the (Becker) coderived category of the exact category $\mathsf{A}$. We refer to [32, Section 7] and [40, Remark 9.2] for a general discussion of the history and philosophy of the coderived and contraderived categories. In the context of the present paper, the discussion in [36, Section 7] is relevant.
Generally speaking, for the abelian category $R{\operatorname{\mathsf{--Mod}}}$ of modules over a ring $R$, the coderived category $\mathsf{D}^{\mathsf{bco}}(R{\operatorname{\mathsf{--Mod}}})$ is quite different from the contraderived category $\mathsf{D}^{\mathsf{bctr}}(R{\operatorname{\mathsf{--Mod}}})$. For the abelian category of $\mathfrak{R}$-contramodules $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$, the construction of the coderived category does not even make sense, as there are no nonzero injective objects in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ in general. However, the coderived category of the exact category of flat $R$-modules agrees with the derived and contraderived categories of the same exact category [36, Sections 7.4 and 7.6]. In this paper we generalize these results to the exact category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ of flat contramodules over a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open two-sided ideals.
For the abelian category of $R$-modules $\mathsf{B}=R{\operatorname{\mathsf{--Mod}}}$, the contraderived category is equivalent to the homotopy category of complexes of projective objects, $\mathsf{Hot}(R{\operatorname{\mathsf{--Mod}}}_{\mathsf{proj}})\simeq\mathsf{D}^{\mathsf{bctr}}(R{\operatorname{\mathsf{--Mod}}})$[26, Proposition 8.1]. The equivalence of two conditions (i) and (ii) above (which is a part of [26, Theorem 8.6]) can be restated by saying that a complex of flat $R$-modules is contraacyclic if and only if it is acyclic as a complex in the exact category of flat $R$-modules $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}$. Besides, it is obvious that a complex in the exact category $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}$ is contraacyclic if and only if it is contraacyclic as a complex in $R{\operatorname{\mathsf{--Mod}}}$, because the projective objects of $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}$ coincide with those of $R{\operatorname{\mathsf{--Mod}}}$. It follows that the contraderived category of $R$-modules is equivalent to the derived category of the exact category of flat $R$-modules, $\mathsf{Hot}(R{\operatorname{\mathsf{--Mod}}}_{\mathsf{proj}})\simeq\mathsf{D}(R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}})=\mathsf{D}^{\mathsf{bctr}}(R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}})\simeq\mathsf{D}^{\mathsf{bctr}}(R{\operatorname{\mathsf{--Mod}}})$.
On the other hand, the injective objects of the exact category $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}$ are the flat cotorsion $R$-modules. So the equivalence of two conditions (ii) and (v) above (which follows immediately from [2, Theorem 5.3]) can be restated by saying that the classes of acyclic and coacyclic complexes in the exact category $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}$ coincide. It also follows from [2, Theorem 5.3] that the coderived category of the exact category $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}$ is equivalent to the homotopy category of complexes of flat cotorsion $R$-modules, $\mathsf{Hot}(R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}^{\mathsf{cot}})\simeq\mathsf{D}^{\mathsf{bco}}(R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}})$. Thus we have $\mathsf{Hot}(R{\operatorname{\mathsf{--Mod}}}_{\mathsf{proj}})\simeq\mathsf{D}^{\mathsf{bctr}}(R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}})=\mathsf{D}(R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}})=\mathsf{D}^{\mathsf{bco}}(R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}})\simeq\mathsf{Hot}(R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}^{\mathsf{cot}})$.
For the abelian category of $\mathfrak{R}$-contramodules $\mathsf{B}=\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$, the contraderived category is also equivalent to the homotopy category of complexes of projective objects, $\mathsf{Hot}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{proj}})\simeq\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$. This is a quite general result, valid for any locally presentable abelian category $\mathsf{B}$ with enough projective objects (which means, in particular, for the category of left contramodules over any complete, separated topological ring $\mathfrak{R}$ with a base of neighborhoods of zero formed by open right ideals) [40, Corollary 7.4]. When $\mathfrak{R}$ has a countable base of neighborhoods of zero consisting of open two-sided ideals, the results of the present paper allow us to say more.
The equivalence of two conditions (ic) and (iic) above (which is claimed in our Theorem 11.1) can be restated by saying that a complex of flat $\mathfrak{R}$-contramodules is contraacyclic if and only if it is acyclic as a complex in the exact category of flat $\mathfrak{R}$-contramodules $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$. Similarly to the module case, it is obvious that a complex in the exact category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ is contraacyclic if and only if it is contraacyclic as a complex in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$, because the projective objects in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ and $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ coincide. It follows that the contraderived category of $\mathfrak{R}$-contramodules is equivalent to the derived category of the exact category of flat contramodules, $\mathsf{Hot}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{proj}})\simeq\mathsf{D}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})=\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})\simeq\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$.
The equivalence of two conditions (iic) and (vic) above (which is another assertion of our Theorem 11.1) can be restated by saying that a complex of flat $\mathfrak{R}$-contramodules is coacyclic in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ if and only if it is acyclic in this exact category. Furthermore, according to our Corollary 9.2, the coderived category of the exact category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ is equivalent to the homotopy category of complexes of flat cotorsion $\mathfrak{R}$-contramodules, $\mathsf{Hot}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}})\simeq\mathsf{D}^{\mathsf{bco}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$. So we have $\mathsf{Hot}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{proj}})\simeq\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})=\mathsf{D}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})=\mathsf{D}^{\mathsf{bco}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})\simeq\mathsf{Hot}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}})$, just as in the module case.
We refer to the paper [43] for a discussion of another generalization of the results of Neeman’s paper [26] and the paper of Bazzoni, Cortés-Izurdiaga, and Estrada [2], from modules to curved DG-modules. In the present paper, we discuss contramodule generalizations of [26, Theorem 8.6] and [2, Theorem 5.3].
Let us explain the connection with the Benson–Goodearl periodity theorem. Let $0\longrightarrow F\longrightarrow P\longrightarrow F\longrightarrow 0$ be a short exact sequence of modules with a flat $R$-module $F$ and a projective $R$-module $P$. Splicing copies of this short exact sequence on top of each other in both the positive and negative cohomological directions, we obtain an unbounded acyclic complex of projective $R$-modules $P^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ with flat $R$-modules of cocycles. So the complex $F^{\text{\smaller\smaller$\scriptstyle\bullet$}}=P^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ satisfies (ii), and consequently it also satisfies (i). Thus the identity morphism of complexes $P^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow F^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ must be homotopic to zero. Hence the complex $P^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is contractible, and it follows that its modules of cocycles are projective. This is the argument of [26, Remark 2.15].
Conversely, given an acyclic complex of projective $R$-modules with flat $R$-modules of cocycles, one can chop it up into short exact sequence pieces, shift them to one and the same cohomological degree, and take the countable direct sum. This produces a short exact sequence of the form $0\longrightarrow F\longrightarrow P\longrightarrow F\longrightarrow 0$. This argument shows that the Benson–Goodearl periodicity theorem is essentially equivalent to the assertion that in any acyclic complex of projective modules with flat modules of cocycles, the modules of cocycles are actually projective [10, proof of Proposition 7.6], [13, Propositions 1 and 2], [2, Proposition 2.4].
Similarly, one can start with the equivalence of two conditions (ic) and (iic) and deduce the following version of flat and projective periodicity theorem for contramodules. If $0\longrightarrow\mathfrak{F}\longrightarrow\mathfrak{P}\longrightarrow\mathfrak{F}\longrightarrow 0$ is a short exact sequence of $\mathfrak{R}$-contramodules with a flat contramodule $\mathfrak{F}$ and a projective contramodule $\mathfrak{P}$, then the contramodule $\mathfrak{F}$ is actually projective. A direct argument deducing this assertion from the Benson–Goodearl periodicity theorem for modules was spelled out in [34, Proposition 12.1]. The equivalence of two conditions (ic) and (iic), proved in Theorem 11.1 in the present paper, is a much stronger result.
Analogous arguments are applicable to the cotorsion periodicity of Bazzoni, Cortés-Izurdiaga, and Estrada. The cotorsion periodicity theorem for $R$-modules, stated above in this introduction, is essentially equivalent to the assertion that in any acyclic complex of cotorsion $R$-modules, the $R$-modules of cocycles are actually cotorsion  [2, Theorem 1.2(2), Proposition 4.8(2), and Theorem 5.1(2)]. Similarly, the cotorsion periodicity theorem for $\mathfrak{R}$-contramodules claims that if $0\longrightarrow\mathfrak{M}\longrightarrow\mathfrak{C}\longrightarrow\mathfrak{M}\longrightarrow 0$ is a short exact sequence of $\mathfrak{R}$-contramodules with a cotorsion contramodule $\mathfrak{C}$, then $\mathfrak{M}$ is also a cotorsion contramodule [34, Theorem 12.3]. This theorem can be restated by saying that in any acyclic complex of cotorsion $\mathfrak{R}$-contramodules $\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$, the contramodules of cocycles are cotorsion [34, Corollary 12.4].
We refer to the introduction to the paper [3] and to the preprint [36, Sections 7.8 and 7.10] for a further discussion of periodicity theorems. Before we finish this introduction, let us say a few words about contramodules.
Contramodules are modules with infinite summation operations. Contramodule categories, dual-analogous to the categories of comodules, or torsion modules, or discrete, or smooth modules, can be assigned to various algebraic structures [30]. In most cases, the category of contramodules is abelian. In particular, to any complete, separated topological ring $\mathfrak{R}$ with a base of neighborhoods of zero formed by open right ideals, one assigns the abelian category of left $\mathfrak{R}$-contramodules $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$.
The categories of topological modules are usually not abelian; and indeed, contramodules are not topological modules. The $\mathfrak{R}$-contramodules are $\mathfrak{R}$-modules endowed with infinite summation operations with the coefficient families converging to zero in the topology of $\mathfrak{R}$[30, Section 2.1], [38, Sections 1.2 and 5], [39, Section 6]. There is a naturally induced topology on a left $\mathfrak{R}$-contramodule, but it need not be even separated. So $\mathfrak{R}$-contramodules are nontopological modules over a topological ring $\mathfrak{R}$, forming an abelian category (in fact, a locally presentable abelian category with enough projective objects; cf. [40, Sections 6–7]).
When the topological ring $\mathfrak{R}$ has a countable base of neighborhoods of zero, the contramodule theory simplifies quite a bit. In particular, all left $\mathfrak{R}$-contramodules are complete (though they still need not be separated) in their induced topologies. Furthermore, there is a well-behaved full subcategory of flat $\mathfrak{R}$-contramodules $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}\subset\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$. Flat left $\mathfrak{R}$-contramodules are separated and complete in their induced topologies. The full subcategory $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ is resolving and closed under directed colimits in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$. Moreover, the directed colimits are exact functors in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$[38, Sections 6–7] (but not in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$[38, Examples 4.4]).
When $\mathfrak{R}$ has a countable base of neighborhoods of zero consisting of open two-sided ideals, flat $\mathfrak{R}$-contramodules can be described as certain projective systems of flat modules over the discrete quotient rings of $\mathfrak{R}$. The latter fact, however, plays almost no role in the present paper, where we study complexes of flat $\mathfrak{R}$-contramodules using $\aleph_{1}$-accessibility and $\aleph_{1}$-presentability results for contramodule categories that were obtained in the paper [34] (which was, in turn, based on the paper [35]). The description of flat $\mathfrak{R}$-contramodules in terms of systems of flat modules over discrete quotient rings of $\mathfrak{R}$ was used in [34].
We also consider the exact category of cotorsion $\mathfrak{R}$-contramodules $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}$. According to [36, Theorem 7.17 and Corollary 7.21], for any associative ring $R$, the inclusion of exact/abelian categories $R{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ (where $R{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}}$ is the full subcategory of cotorsion left $R$-modules) induces equivalences of their derived and contraderived categories $\mathsf{D}(R{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}})\simeq\mathsf{D}(R{\operatorname{\mathsf{--Mod}}})$ and $\mathsf{D}^{\mathsf{bctr}}(R{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}})\simeq\mathsf{D}^{\mathsf{bctr}}(R{\operatorname{\mathsf{--Mod}}})$. For a complete, separated topological ring $\mathfrak{R}$ with a countable base of neighborhoods of zero consisting of two-sided ideals, it was shown in [34, Corollary 12.8] that the inclusion of exact/abelian categories $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ induces an equivalence of the derived categories $\mathsf{D}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})\simeq\mathsf{D}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$. In this paper we show that the same inclusion induces an equivalence of the contraderived categories $\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})\simeq\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$.
Geometrically, flat contramodules over a commutative topological ring with a countable topology base of open ideals can be identified with flat pro-quasi-coherent pro-sheaves on an ind-affine $\aleph_{0}$-ind-scheme [5, Section 7.11.3], [31, Examples 3.8], [34, Example 8.9]. This constitutes a geometric motivation for the present work.
This paper consists roughly of two parts. Sections 14 present the preliminary material, and then in Sections 56 we prove contramodule versions of results related to the flat/projective periodicity theorem of Benson–Goodearl and Neeman. Section 7 spells out additional preliminaries for the second part, and then in Sections 810 contramodule versions of results related to the cotorsion periodicity theorem of Bazzoni, Cortés-Izurdiaga, and Estrada are proved. The final Section 11 summarizes the results obtained in Sections 56 and 810.
Acknowledgement
I am grateful to Jan Št’ovíček for helpful discussions. The author is supported by the GAČR project 23-05148S and the Czech Academy of Sciences (RVO 67985840).
1. Accessible Subcategories
We use the book [1] as a general reference source on accessible and locally presentable categories. In particular, we refer to [1, Definition 1.4, Theorem 1.5, Corollary 1.5, Definition 1.13(1), and Remark 1.21] for a discussion of $\kappa$-directed colimits (indexed by $\kappa$-directed posets) vs. $\kappa$-filtered colimits (indexed by $\kappa$-filtered small categories). Here $\kappa$ is a regular cardinal.
Let $\mathsf{K}$ be a category with $\kappa$-directed (equivalently, $\kappa$-filtered) colimits. An object $S\in\mathsf{K}$ is said to be $\kappa$-presentable if the functor $\operatorname{Hom}_{\mathsf{K}}(S,{-})\colon\mathsf{K}\longrightarrow\mathsf{Sets}$ from $\mathsf{K}$ to the category of sets $\mathsf{Sets}$ preserves $\kappa$-directed colimits. When the category $\mathsf{K}$ is (pre)additive, an object $S\in\mathsf{K}$ is $\kappa$-presentable if and only if the functor $\operatorname{Hom}_{\mathsf{K}}(S,{-})\colon\mathsf{K}\longrightarrow\mathsf{Ab}$ from $\mathsf{K}$ to the category of abelian groups $\mathsf{Ab}$ preserves $\kappa$-directed colimits.
The category $\mathsf{K}$ is said to be $\kappa$-accessible [1, Definition 2.1] if there exists a set of $\kappa$-presentable objects $\mathsf{S}\subset\mathsf{K}$ such that every object of $\mathsf{K}$ is a $\kappa$-directed colimit of objects from $\mathsf{S}$. If this is the case, then the $\kappa$-presentable objects of $\mathsf{K}$ are precisely all the retracts (direct summands) of the objects from $\mathsf{S}$. The category $\mathsf{K}$ is said to be locally $\kappa$-presentable [1, Definition 1.17 and Theorem 1.20] if $\mathsf{K}$ is $\kappa$-accessible and all colimits exists in $\mathsf{K}$.
In the case of the countable cardinal $\kappa=\aleph_{0}$, the $\aleph_{0}$-presentable objects are known as finitely presentable [1, Definition 1.1], the $\aleph_{0}$-accessible categories are called finitely accessible [1, Remark 2.2(1)], and the locally $\aleph_{0}$-presentable categories are known as locally finitely presentable [1, Definition 1.9 and Theorem 1.11]. In the case of the cardinal $\kappa=\aleph_{1}$, we will call the $\aleph_{1}$-presentable objects countably presentable.
[…]
Proposition 1.1.
[…]
Proof.
This well-known result goes back, at least, to [24, Proposition 2.1], [11, Section 4.1], and [23, Proposition 5.11] (notice that what we call finitely accessible categories were called “locally finitely presented categories” in [11, 23]). The nontrivial part is the “if” implication in the last assertion; all the other assertions follow easily. See [35, Proposition 1.2] for some details. ∎
Let $\alpha$ be an ordinal and $\mathsf{K}$ be a category. An $\alpha$-indexed chain (of objects and morphisms) in $\mathsf{K}$ is a commutative diagram $(K_{i}\to K_{j})_{0\leq i<j<\alpha}$ in $\mathsf{K}$ indexed by the directed poset $\alpha$.
Given an additive/abelian category $\mathsf{K}$, we denote by $\mathsf{Com}(\mathsf{K})$ the additive/abelian category of (unbounded) complexes in $\mathsf{K}$.
Proposition 1.2.
Let $\kappa$ be an uncountable regular cardinal and $\lambda<\kappa$ be a smaller infinite cardinal.
(a) Let $\mathsf{K}$ be an additive category with $\kappa$-directed colimits. Then all complexes of $\kappa$-presentable objects in $\mathsf{K}$ are $\kappa$-presentable as objects of the category of complexes $\mathsf{Com}(\mathsf{K})$.
(b) Let $\mathsf{K}$ be a $\kappa$-accessible additive category where the colimits of $\lambda$-indexed chains exist. Then the category $\mathsf{Com}(\mathsf{K})$ of complexes in $\mathsf{K}$ is $\kappa$-accessible. The $\kappa$-presentable objects of $\mathsf{Com}(\mathsf{K})$ are precisely all the complexes of $\kappa$-presentable objects in $\mathsf{K}$.
Proof.
Part (a) holds because $\kappa$-directed colimits commute with countable limits in the category $\mathsf{Sets}$ or $\mathsf{Ab}$, cf. [34, Lemma 1.1]. Part (b) is a particular case of [35, Theorem 6.2]; cf. [35, proof of Corollary 10.4]. ∎
2. Contraderived Categories
We refer to the survey paper [8] for general background on exact categories in the sense of Quillen. The Yoneda Ext groups in an exact category $\mathsf{K}$ can be defined, e. g., as $\operatorname{Ext}^{n}_{\mathsf{K}}(X,Y)=\operatorname{Hom}_{\mathsf{D}(\mathsf{K})}(X,Y[n])$ for all $X$, $Y\in\mathsf{K}$ and $n\geq 0$. Here $\mathsf{D}(\mathsf{K})$ denotes the derived category of $\mathsf{K}$.
Let $\mathsf{K}$ be an exact category, and let $\mathsf{F}$, $\mathsf{C}\subset\mathsf{K}$ be two classes of objects. One denotes by $\mathsf{F}^{\perp_{1}}\subset\mathsf{K}$ the class of all objects $X\in\mathsf{K}$ such that $\operatorname{Ext}^{1}_{\mathsf{K}}(F,X)=0$ for all $F\in\mathsf{F}$. Dually, the notation ${}^{\perp_{1}}\mathsf{C}\subset\mathsf{K}$ stands for the class of all objects $Y\in\mathsf{K}$ such that $\operatorname{Ext}^{1}_{\mathsf{K}}(Y,C)=0$ for all $C\in\mathsf{C}$. Similarly, $\mathsf{F}^{\perp_{\geq 1}}\subset\mathsf{K}$ is the class of all objects $X\in\mathsf{K}$ such that $\operatorname{Ext}^{n}_{\mathsf{K}}(F,X)=0$ for all $n\geq 1$ and $F\in\mathsf{F}$, while ${}^{\perp_{\geq 1}}\mathsf{C}\subset\mathsf{K}$ is the class of all objects $Y\in\mathsf{K}$ such that $\operatorname{Ext}^{n}_{\mathsf{K}}(Y,C)=0$ for all $n\geq 1$ and $C\in\mathsf{C}$.
Given an exact category $\mathsf{K}$, any full subcategory $\mathsf{E}\subset\mathsf{K}$ closed under extensions in $\mathsf{K}$ can be endowed with an exact category structure in which the (admissible) short exact sequences in $\mathsf{E}$ are the short exact sequences in $\mathsf{K}$ with the terms belonging to $\mathsf{E}$. We will say that the exact structure on $\mathsf{E}$ is inherited from the exact structure on $\mathsf{K}$.
A full subcategory $\mathsf{F}\subset\mathsf{K}$ is said to be self-generating if for every admissible epimorphism $K\longrightarrow F$ in $\mathsf{K}$ with $F\in\mathsf{F}$ there exists a morphism $G\longrightarrow K$ in $\mathsf{K}$ with $G\in\mathsf{F}$ such that the composition $G\longrightarrow K\longrightarrow F$ is an admissible epimorphism $G\longrightarrow F$ in $\mathsf{K}$. A full subcategory $\mathsf{F}\subset\mathsf{K}$ is said to be generating if for every object $K\in\mathsf{K}$ there exists an object $G\in\mathsf{F}$ together with an admissible epimorphism $G\longrightarrow K$ in $\mathsf{K}$. Clearly, any generating full subcategory is self-generating.
A full subcategory $\mathsf{F}\subset\mathsf{K}$ is said to be self-resolving if it is self-generating and closed under extensions and kernels of admissible epimorphisms. We will say that a full subcategory $\mathsf{F}\subset\mathsf{K}$ is resolving if it is generating and closed under extensions and kernels of admissible epimorphisms. Clearly, any resolving full subcategory is self-resolving.
Lemma 2.1.
Let $\mathsf{K}$ be an exact category and $\mathsf{F}\subset\mathsf{K}$ be a self-generating full subcategory closed under kernels of admissible epimorphisms. Then one has $\mathsf{F}^{\perp_{1}}=\mathsf{F}^{\perp_{\geq 1}}\subset\mathsf{K}$.
Proof.
The argument from [46, Lemma 6.17] applies. See [3, Lemma 1.3] for some additional details. ∎
[…]
Given a class of objects $\mathsf{S}\subset\mathsf{K}$, we denote by $\mathsf{Fil}(\mathsf{S})\subset\mathsf{K}$ the class of all objects filtered by (objects isomorphic to) the objects from $\mathsf{S}$. A class of objects $\mathsf{F}\subset\mathsf{K}$ is said to be deconstructible if there exists a set of objects $\mathsf{S}\subset\mathsf{K}$ such that $\mathsf{F}=\mathsf{Fil}(\mathsf{S})$. The following result is known classically as the Eklof lemma [12, Lemma 1].
Lemma 2.2.
Let $\mathsf{K}$ be an abelian category with infinite coproducts. Then, for any class of objects $\mathsf{C}\subset\mathsf{K}$, the class of objects ${}^{\perp_{1}}\mathsf{C}\subset\mathsf{K}$ is closed under transfinitely iterated extensions in $\mathsf{K}$. In other words, $\mathsf{Fil}({}^{\perp_{1}}\mathsf{C})\subset{}^{\perp_{1}}\mathsf{C}$.
Proof.
In the stated generality, a proof can be found in [38, Lemma 4.5]; see also [40, Proposition 1.3] and [42, Lemma 7.5]. ∎
The following lemma is well known and easy.
Lemma 2.3.
Let $\mathsf{B}$ be an abelian category, $A\in\mathsf{B}$ be an object, and $n\in\mathbb{Z}$ be an integer. Denote by $D_{n,n+1}^{\text{\smaller\smaller$\scriptstyle\bullet$}}(A)$ the contractible two-term complex $\dotsb\longrightarrow 0\longrightarrow A\overset{{\mathrm{id}}}{\longrightarrow}A\longrightarrow 0\longrightarrow\dotsb$ concentrated in the cohomological degrees $n$ and $n+1$. Then, for any complex $B^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ in $\mathsf{B}$ and all integers $i\geq 0$ there is a natural isomorphism of abelian groups
$\operatorname{Ext}^{i}_{\mathsf{Com}(\mathsf{B})}(D_{n,n+1}^{\text{\smaller\smaller$\scriptstyle\bullet$}}(A),B^{\text{\smaller\smaller$\scriptstyle\bullet$}})\simeq\operatorname{Ext}^{i}_{\mathsf{B}}(A,B^{n}).$
Proof.
See, e. g., [15, Lemma 3.1(5)] or [33, Lemma 1.8]. ∎
Given an additive category $\mathsf{B}$, we denote by $\mathsf{Hot}(\mathsf{B})$ the triangulated category of (unbounded) complexes in $\mathsf{B}$ with morphisms up to cochain homotopy. The notation $B^{\text{\smaller\smaller$\scriptstyle\bullet$}}[i]$,  $i\in\mathbb{Z}$ stands for the cohomological degree shift of a complex $B^{\text{\smaller\smaller$\scriptstyle\bullet$}}$; so $(B^{\text{\smaller\smaller$\scriptstyle\bullet$}}[i])^{n}=B^{n+i}$ for all $n\in\mathbb{Z}$. The following lemma is well known.
Lemma 2.4.
Let $\mathsf{B}$ be an abelian category, and let $A^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ and $B^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be two complexes in $\mathsf{B}$. Assume that $\operatorname{Ext}^{1}_{\mathsf{B}}(A^{n},B^{n})=0$ for all $n\in\mathbb{Z}$. Then there is a natural isomorphism of abelian groups
$\operatorname{Ext}^{1}_{\mathsf{Com}(\mathsf{B})}(A^{\text{\smaller\smaller$\scriptstyle\bullet$}},B^{\text{\smaller\smaller$\scriptstyle\bullet$}})\simeq\operatorname{Hom}_{\mathsf{Hot}(\mathsf{B})}(A^{\text{\smaller\smaller$\scriptstyle\bullet$}},B^{\text{\smaller\smaller$\scriptstyle\bullet$}}[1]).$
Proof.
See, e. g., [15, Lemma 2.1], [3, Lemma 1.6] or [33, Lemma 1.9]. ∎
Let $\mathsf{B}$ be an abelian (or exact) category with enough projective objects. A complex $B^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ in $\mathsf{B}$ is said to be contraacyclic (in the sense of Becker [4, Proposition 1.3.6(1)]) if, for every complex of projective objects $P^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ in $\mathsf{B}$, the complex of abelian groups $\operatorname{Hom}_{\mathsf{B}}^{\text{\smaller\smaller$\scriptstyle\bullet$}}(P^{\text{\smaller\smaller$\scriptstyle\bullet$}},B^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ is acyclic. The full subcategory of contraacyclic complexes is denoted by $\mathsf{Ac}^{\mathsf{bctr}}(\mathsf{B})\subset\mathsf{Hot}(\mathsf{B})$ or (depending on context) $\mathsf{Ac}^{\mathsf{bctr}}(\mathsf{B})\subset\mathsf{Com}(\mathsf{B})$. The triangulated Verdier quotient category
$\mathsf{D}^{\mathsf{bctr}}(\mathsf{B})=\mathsf{Hot}(\mathsf{B})/\mathsf{Ac}^{\mathsf{bctr}}(\mathsf{B})$
is called the (Becker) contraderived category of $\mathsf{B}$.
We are interested in (termwise admissible) short exact sequences $0\longrightarrow K^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow L^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow M^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow 0$ of complexes in $\mathsf{B}$. Any such short exact sequence can be viewed as a bicomplex with three rows in $\mathsf{B}$. Then one can consider its total complex, which we will denote by $\operatorname{Tot}(K^{\text{\smaller\smaller$\scriptstyle\bullet$}}\to L^{\text{\smaller\smaller$\scriptstyle\bullet$}}\to M^{\text{\smaller\smaller$\scriptstyle\bullet$}})$.
Lemma 2.5.
Let $\mathsf{B}$ be an exact category with enough projective objects. Then
(a) for any short exact sequence $0\longrightarrow K^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow L^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow M^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow 0$ of complexes in $\mathsf{B}$, the total complex $\operatorname{Tot}(K^{\text{\smaller\smaller$\scriptstyle\bullet$}}\to L^{\text{\smaller\smaller$\scriptstyle\bullet$}}\to M^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ is contraacyclic;
(b) assuming that infinite products exists in $\mathsf{B}$, the class of all contraacyclic complexes is closed under infinite products.
Proof.
These observations seem to go back to [28, Section 3.5] and [22, Proposition 4.3]. Part (b) is obvious. In part (a), the point is that, for every complex of projective objects $P^{\text{\smaller\smaller$\scriptstyle\bullet$}}\in\mathsf{B}$, the short sequence of complexes of abelian groups $0\longrightarrow\operatorname{Hom}_{\mathsf{B}}^{\text{\smaller\smaller$\scriptstyle\bullet$}}(P^{\text{\smaller\smaller$\scriptstyle\bullet$}},K^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow\operatorname{Hom}_{\mathsf{B}}^{\text{\smaller\smaller$\scriptstyle\bullet$}}(P^{\text{\smaller\smaller$\scriptstyle\bullet$}},L^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow\operatorname{Hom}_{\mathsf{B}}^{\text{\smaller\smaller$\scriptstyle\bullet$}}(P^{\text{\smaller\smaller$\scriptstyle\bullet$}},M^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow 0$ is exact. The total complex of any short exact sequence of complexes of abelian groups is exact. For a generalization, see Lemma 7.5 below. ∎
Lemma 2.6.
Let $\mathsf{B}$ be an abelian category with enough projective objects. Then any contraacyclic complex in $\mathsf{B}$ is acyclic.
Proof.
This is the dual assertion to [31, Lemma A.2] or a particular case of [29, Lemma B.7.3(b)]. It is not known whether the assertion of this lemma holds for exact categories in general; see [29, Remark B.7.4] for a discussion. ∎
A complex in an exact category $\mathsf{E}$ is said to be absolutely acyclic [27, Section 2.1], [28, Section 3.3], [29, Appendix A] if it belongs to the minimal thick subcategory of $\mathsf{Hot}(\mathsf{E})$ containing the total complexes of (termwise admissible) short exact sequences of complexes in $\mathsf{E}$. Equivalently, a complex in $\mathsf{E}$ is absolutely acyclic if and only if it belongs to the minimal full subcategory of $\mathsf{Com}(\mathsf{E})$ containing contractible complexes and closed under extensions and direct summands [41, Proposition 8.12].
It is clear from Lemma 2.5(a) that all absolutely acyclic complexes in $\mathsf{B}$ are contraacyclic.
An exact category $\mathsf{E}$ is said to have homological dimension $\leq\nobreak d$ (where $d\geq-1$ is an integer) if $\operatorname{Ext}^{d+1}_{\mathsf{E}}(X,Y)=0$ for all $X$, $Y\in\mathsf{E}$.
Lemma 2.7.
In an exact category of finite homological dimension, all acyclic complexes are absolutely acyclic.
Proof.
This is [27, Remark 2.1]. For additional information, see [34, Proposition 6.2]. ∎
Let us denote by $\mathsf{B}_{\mathsf{proj}}\subset\mathsf{B}$ the full subcategory of projective objects in an abelian/exact category $\mathsf{B}$.
Theorem 2.8.
For any locally presentable abelian category $\mathsf{B}$ with enough projective objects, the inclusion of additive/abelian categories $\mathsf{B}_{\mathsf{proj}}\longrightarrow\mathsf{B}$ induces a triangulated equivalence
$\mathsf{Hot}(\mathsf{B}_{\mathsf{proj}})\simeq\mathsf{D}^{\mathsf{bctr}}(\mathsf{B}).$
Proof.
This is [40, Corollary 7.4]. A far-reaching generalization can be found in [41, Corollary 6.14]. See also Theorem 7.2 below. ∎
Our final lemma in this section is standard and easy, but helpful to keep in mind.
Lemma 2.9.
Let $\mathsf{H}$ be a triangulated category, $\mathsf{X}\subset\mathsf{H}$ be a thick subcategory, and $\mathsf{F}\subset\mathsf{H}$ be a full triangulated subcategory. Assume that for every object $H\in\mathsf{H}$ there exists an object $F\in\mathsf{F}$ together with a morphism $F\longrightarrow H$ whose cone belongs to $\mathsf{X}$. Then the inclusion functor $\mathsf{F}\longrightarrow\mathsf{H}$ induces a triangulated equivalence of the Verdier quotient categories
$\mathsf{F}/(\mathsf{X}\cap\mathsf{F})\simeq\mathsf{H}/\mathsf{X}.$
Proof.
This is [20, Proposition 10.2.7(ii)] or [28, Lemma 1.6(a)]. ∎
3. Topological Rings
The definition of a contramodule over a topological ring goes back to [27, Remark A.3]. We refer to [30, Section 2.1], [38, Sections 1.2 and 5], and [34, Sections 7–8] for further details and references in connection with the definitions and assertions presented below in this section. The pedagogical exposition in [39, Sections 6–7] is particularly recommended.
[…]
Given an abelian group $A$ and a set $X$, we denote by $A[X]=A^{(X)}=\bigoplus_{x\in X}A$ the direct sum of copies of the group $A$ indexed by the set $X$. The elements of $A[X]$ are interpreted as finite formal linear combitations $\sum_{x\in X}a_{x}x$ of elements of $X$ with the coefficients $a_{x}\in A$. So one has $a_{x}=0$ for all but a finite subset of indices $x$.
[…]
For any map of sets $f\colon X\longrightarrow Y$, the induced map $\mathfrak{A}[[f]]\colon\mathfrak{A}[[X]]\longrightarrow\mathfrak{A}[[Y]]$ takes a formal linear combination $\sum_{x\in X}a_{x}x$ to the formal linear combination $\sum_{y\in Y}b_{y}y$ with the coefficients defined by the rule $b_{y}=\sum_{x\in X}^{f(x)=y}a_{x}$. Here the infinite summation sign in the latter formula denotes the limit of finite partial sums in the topology of $\mathfrak{A}$. So the assignment $X\longmapsto\mathfrak{A}[[X]]$ is a functor $\mathsf{Sets}\longrightarrow\mathsf{Ab}$, but we will mostly consider it as a functor $\mathsf{Sets}\longrightarrow\mathsf{Sets}$.
Now let $\mathfrak{R}$ be a complete, separated topological ring with a base of neighborhoods of zero formed by open right ideals. Then the functor $\mathfrak{R}[[{-}]]\colon\mathsf{Sets}\longrightarrow\mathsf{Sets}$ has a natural structure of a monad on the category of sets (in the sense of [25, Chapter VI]). For any set $X$, the monad unit $\epsilon_{\mathfrak{R},X}\colon X\longrightarrow\mathfrak{R}[[X]]$ is the map taking every element $x\in X$ to the formal linear combination $\sum_{y\in X}r_{y}y$ with $r_{x}=1$ and $r_{y}=0$ for $y\neq x$. The monad multiplication $\phi_{\mathfrak{R},X}\colon\mathfrak{R}[[\mathfrak{R}[[X]]]]\longrightarrow\mathfrak{R}[[X]]$ is the “opening of parentheses” map assigning a formal linear combination to a formal linear combination of formal linear combinations. In order to open the parentheses, one needs to compute pairwise products of elements of the ring $\mathfrak{R}$ and certain infinite sums of elements of $\mathfrak{R}$, which are understood as the limits of finite partial sums in the topology of $\mathfrak{R}$. The conditions imposed on $\mathfrak{R}$ (including, in particular, the requirement of a topology base of open right ideals) guarantee the convergence.
In the general category-theoretic terminology, one usually speaks about “algebras over monads”, but in our context we prefer to call them modules over monads. A left $\mathfrak{R}$-contramodule is defined as a module over the monad $X\longmapsto\mathfrak{R}[[X]]$ on the category of sets. In other words, a left $\mathfrak{R}$-contramodule $\mathfrak{P}$ is a set endowed with a left contraaction map $\pi_{\mathfrak{P}}\colon\mathfrak{R}[[\mathfrak{P}]]\longrightarrow\mathfrak{P}$ satisfying the following contraassociativity and contraunitality axioms. Firstly, the “opening of parentheses” map $\phi_{\mathfrak{R},\mathfrak{P}}\colon\mathfrak{R}[[\mathfrak{R}[[\mathfrak{P}]]]]\longrightarrow\mathfrak{R}[[\mathfrak{P}]]$ and the map $\mathfrak{R}[[\pi_{\mathfrak{P}}]]\colon\mathfrak{R}[[\mathfrak{R}[[\mathfrak{P}]]]]\longrightarrow\mathfrak{R}[[\mathfrak{P}]]$ induced by the map $\pi_{\mathfrak{P}}$ must have equal compositions with the contraaction map $\pi_{\mathfrak{P}}$,
$\mathfrak{R}[[\mathfrak{R}[[\mathfrak{P}]]]]\rightrightarrows\mathfrak{R}[[\mathfrak{P}]]\longrightarrow\mathfrak{P}.$
Secondly, the composition of the map $\epsilon_{\mathfrak{R},\mathfrak{P}}\colon\mathfrak{P}\longrightarrow\mathfrak{R}[[\mathfrak{P}]]$ with the map $\pi_{\mathfrak{P}}\colon\mathfrak{R}[[\mathfrak{P}]]\allowbreak\longrightarrow\mathfrak{P}$ must be equal to the identity map ${\mathrm{id}}_{\mathfrak{P}}$,
[…]
In particular, given a discrete ring $R$, the similar (but simpler) construction produces a monad structure on the functor $X\longmapsto R[X]\colon\mathsf{Sets}\longrightarrow\mathsf{Sets}$. Modules over this monad on $\mathsf{Sets}$ are the same things as the usual left $R$-modules. Now we have two monads on $\mathsf{Sets}$ associated with a topological ring $\mathfrak{R}$ satisfying the conditions above: the monad $X\longmapsto\mathfrak{R}[[X]]$ (depending on the topology on $\mathfrak{R}$) and the monad $X\longmapsto\mathfrak{R}[X]$ (defined irrespectively of any topology on $\mathfrak{R}$). The inclusion of the set of all finite formal linear combinations into the set of zero-convergent infinite ones is a natural injective map $\mathfrak{R}[X]\longrightarrow\mathfrak{R}[[X]]$. Given a left $\mathfrak{R}$-contramodule structure on a set $\mathfrak{P}$, the composition $\mathfrak{R}[\mathfrak{P}]\longrightarrow\mathfrak{R}[[\mathfrak{P}]]\longrightarrow\mathfrak{P}$ endows $\mathfrak{P}$ with its underlying left $\mathfrak{R}$-module structure. In other words, the finite aspects of the contramodule infinite summation operations define a module structure.
The category of left $\mathfrak{R}$-contramodules $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ is abelian with exact functors of infinite products. The forgetful functor $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Mod}}}$ from the category of left $\mathfrak{R}$-contramodules to the category of left $\mathfrak{R}$-modules is exact and preserves infinite products. We will denote the $\operatorname{Hom}$ and $\operatorname{Ext}$ groups computed in the abelian category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ by $\operatorname{Hom}^{\mathfrak{R}}(\mathfrak{P},\mathfrak{Q})$ and $\operatorname{Ext}^{\mathfrak{R},*}(\mathfrak{P},\mathfrak{Q})$.
Furthermore, for any set $X$ the contraaction map $\pi_{\mathfrak{R}[[X]]}=\phi_{\mathfrak{R},X}\colon\mathfrak{R}[[\mathfrak{R}[[X]]]]\longrightarrow\mathfrak{R}[[X]]$ defines a left $\mathfrak{R}$-contramodule structure on the set $\mathfrak{R}[[X]]$. The resulting left $\mathfrak{R}$-contramodule $\mathfrak{R}[[X]]$ is called the free $\mathfrak{R}$-contramodule spanned by the set $X$. For any left $\mathfrak{R}$-contramodule $\mathfrak{P}$, the group of all $\mathfrak{R}$-contramodule morphisms $\mathfrak{R}[[X]]\longrightarrow\mathfrak{P}$ is naturally bijective to the group of all maps $X\longrightarrow\mathfrak{P}$,
$\operatorname{Hom}^{\mathfrak{R}}(\mathfrak{R}[[X]],\mathfrak{P})\simeq\operatorname{Hom}_{\mathsf{Sets}}(X,\mathfrak{P}).$
It follows easily that there are enough projective objects in the abelian category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$. The projective objects of $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ are precisely all the direct summands of free $\mathfrak{R}$-contramodules. In particular, the free left $\mathfrak{R}$-contramodule with one generator $\mathfrak{R}=\mathfrak{R}[[\{*\}]]$ is a projective generator of $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$. The abelian category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ is locally $\kappa$-presentable, where $\kappa$ is any cardinal greater than the cardinality of a base of neighborhoods of zero in $\mathfrak{R}$.
A right $\mathfrak{R}$-module $\mathcal{N}$ is said to be discrete if the action map $\mathcal{N}\times\mathfrak{R}\longrightarrow\mathcal{N}$ is continuous in the given topology of $\mathfrak{R}$ and the discrete topology of $\mathcal{N}$. The full subcategory of discrete right $\mathfrak{R}$-modules ${\operatorname{\mathsf{Discr--}}}\mathfrak{R}\subset{\operatorname{\mathsf{Mod--}}}\mathfrak{R}$ is closed under subobjects, quotients, and infinite direct sums in the category of right $\mathfrak{R}$-modules ${\operatorname{\mathsf{Mod--}}}\mathfrak{R}$. The category ${\operatorname{\mathsf{Discr--}}}\mathfrak{R}$ is a Grothendieck abelian category.
The contratensor product $\mathcal{N}\odot_{\mathfrak{R}}\mathfrak{P}$ of a discrete right $\mathfrak{R}$-module $\mathcal{N}$ and a left $\mathfrak{R}$-contramodule $\mathfrak{P}$ is an abelian group constructed as the cokernel of (the difference of) the natural pair of maps
$\mathcal{N}\otimes_{\mathbb{Z}}\mathfrak{R}[[\mathfrak{P}]]\rightrightarrows\mathcal{N}\otimes_{\mathbb{Z}}\mathfrak{P}.$
In this pair of parallel abelian group maps, the first map $\mathcal{N}\otimes_{\mathbb{Z}}\mathfrak{R}[[\mathfrak{P}]]\longrightarrow\mathcal{N}\otimes_{\mathbb{Z}}\mathfrak{P}$ is the map $\mathcal{N}\otimes_{\mathbb{Z}}\pi_{\mathfrak{P}}$ induced by the contraaction map $\pi_{\mathfrak{P}}\colon\mathfrak{R}[[\mathfrak{P}]]\longrightarrow\mathfrak{P}$. The second map $\mathcal{N}\otimes_{\mathbb{Z}}\mathfrak{R}[[\mathfrak{P}]]\longrightarrow\mathcal{N}\otimes_{\mathbb{Z}}\mathfrak{P}$ is the composition $\mathcal{N}\otimes_{\mathbb{Z}}\mathfrak{R}[[\mathfrak{P}]]\longrightarrow\mathcal{N}[\mathfrak{P}]\longrightarrow\mathcal{N}\otimes_{\mathbb{Z}}\mathfrak{P}$ of the map $\mathcal{N}\otimes_{\mathbb{Z}}\mathfrak{R}[[\mathfrak{P}]]\longrightarrow\mathcal{N}[\mathfrak{P}]$ induced by the discrete right action of $\mathfrak{R}$ in $\mathcal{N}$ with a natural surjective map $\mathcal{N}[\mathfrak{P}]\longrightarrow\mathcal{N}\otimes_{\mathbb{Z}}\mathfrak{P}$. Here the map $\mathcal{N}\otimes_{\mathbb{Z}}\mathfrak{R}[[\mathfrak{P}]]\longrightarrow\mathcal{N}[\mathfrak{P}]$ is defined by the rule $n\otimes\sum_{p\in\mathfrak{P}}r_{p}p\longmapsto\sum_{p\in\mathfrak{P}}(nr_{p})p$ for all $n\in\mathcal{N}$ and $\sum_{p\in\mathfrak{P}}r_{p}p\in\mathfrak{R}[[\mathfrak{P}]]$. The sum in the right-hand side is finite, because the family of elements $r_{p}$ is zero-convergent in $\mathfrak{R}$, while the right action of $\mathfrak{R}$ in $\mathcal{N}$ is discrete; so $nr_{p}=0$ in $\mathcal{N}$ for all but a finite subset of elements $p\in\mathfrak{P}$. A natural surjective map $A[B]\longrightarrow A\otimes_{\mathbb{Z}}B$ is defined in the obvious way for any abelian groups $A$ and $B$.
The contratensor product functor $\odot_{\mathfrak{R}}\colon{\operatorname{\mathsf{Discr--}}}\mathfrak{R}\times\mathfrak{R}{\operatorname{\mathsf{--Contra}}}\longrightarrow\mathsf{Ab}$ is right exact and preserves coproducts (in other words, preserves all colimits) in both of its arguments. For any set $X$ and any discrete right $\mathfrak{R}$-module $\mathcal{N}$, there is a natural isomorphism of abelian groups
$\mathcal{N}\odot_{\mathfrak{R}}\mathfrak{R}[[X]]\simeq\mathcal{N}[X]=\mathcal{N}^{(X)}.$
A left $\mathfrak{R}$-contramodule $\mathfrak{F}$ is called flat if the contratensor product functor ${-}\odot_{\mathfrak{R}}\nobreak\mathfrak{F}\colon\allowbreak{\operatorname{\mathsf{Discr--}}}\mathfrak{R}\longrightarrow\mathsf{Ab}$ is exact. It follows that the class of all flat $\mathfrak{R}$-contramodules is preserved by the directed colimits in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ (because the directed colimits are exact in $\mathsf{Ab}$), and all the projective $\mathfrak{R}$-contramodules are flat. We will denote the full subcategory of flat left $\mathfrak{R}$-contramodules by $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}\subset\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$.
Given a closed subgroup $\mathfrak{A}\subset\mathfrak{R}$ and a left $\mathfrak{R}$-contramodule $\mathfrak{P}$, we denote by $\mathfrak{A}\rightthreetimes\mathfrak{P}\subset\mathfrak{P}$ the image of the composition $\mathfrak{A}[[\mathfrak{P}]]\longrightarrow\mathfrak{R}[[\mathfrak{P}]]\longrightarrow\mathfrak{P}$ of the obvious inclusion $\mathfrak{A}[[\mathfrak{P}]]\hookrightarrow\mathfrak{R}[[\mathfrak{P}]]$ with the contraaction map $\pi_{\mathfrak{P}}\colon\mathfrak{R}[[\mathfrak{P}]]\longrightarrow\mathfrak{P}$. For any closed left ideal $\mathfrak{J}\subset\mathfrak{R}$ and any left $\mathfrak{R}$-contramodule $\mathfrak{P}$, the subgroup $\mathfrak{J}\rightthreetimes\mathfrak{P}$ is a subcontramodule in $\mathfrak{P}$. For any open right ideal $\mathfrak{I}\subset\mathfrak{R}$ and any left $\mathfrak{R}$-contramodule $\mathfrak{P}$, there is a natural isomorphism of abelian groups
$(\mathfrak{R}/\mathfrak{I})\odot_{\mathfrak{R}}\mathfrak{P}\simeq\mathfrak{P}/(\mathfrak{I}\rightthreetimes\mathfrak{P}).$
For any closed right ideal $\mathfrak{J}\subset\mathfrak{R}$ and any set $X$, one has
$\mathfrak{J}\rightthreetimes\mathfrak{R}[[X]]=\mathfrak{J}[[X]]\subset\mathfrak{R}[[X]].$
In particular, for an open right ideal $\mathfrak{I}\subset\mathfrak{R}$, one has
$(\mathfrak{R}/\mathfrak{I})\odot_{\mathfrak{R}}\mathfrak{R}[[X]]\simeq\mathfrak{R}[[X]]/\mathfrak{I}[[X]]\simeq(\mathfrak{R}/\mathfrak{I})[X].$
[…]
4. Countably Presentable Contramodules
The exposition in this section is based on [38, Section 6] and [34, Sections 8–10]. In the case of a topological ring with a base of neighborhoods of zero formed by two-sided ideals, [29, Section E.1] is a relevant reference.
Let $\mathfrak{R}$ be a complete, separated topological ring with a base of neighborhoods of zero formed by open right ideals. Given a left $\mathfrak{R}$-contramodule $\mathfrak{P}$, consider the natural map
[…]
Here $\mathfrak{I}$ ranges over all the open right ideals in $\mathfrak{R}$. A contramodule $\mathfrak{P}$ is called separated if the map $\lambda_{\mathfrak{R},\mathfrak{P}}$ is injective (equivalently, this means that $\bigcap_{\mathfrak{I}\subset\mathfrak{R}}(\mathfrak{I}\rightthreetimes\mathfrak{P})=0$ in $\mathfrak{P}$). A contramodule $\mathfrak{P}$ is called complete if the map $\lambda_{\mathfrak{R},\mathfrak{P}}$ is surjective.
Lemma 4.1.
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open right ideals. Then
(a) all left $\mathfrak{R}$-contramodules are complete (but not necessarily separated);
(b) all flat left $\mathfrak{R}$-contramodules are separated.
Proof.
This is [34, Lemmas 8.1 and 8.3]. ∎
Lemma 4.2.
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open right ideals. Then the full subcategory of flat left $\mathfrak{R}$-contramodules $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ is resolving in the abelian category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$.
Proof.
This is [34, Lemma 8.4(a)]. ∎
In particular, it follows from Lemma 4.2 that the full subcategory $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ inherits an exact category structure from the abelian exact structure on $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$. We will consider $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ as an exact category with this exact category structure. Recall that the full subcategory $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ is preserved by the directed colimits in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$.
Lemma 4.3.
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open right ideals. In this context:
(a) For any discrete right $\mathfrak{R}$-module $\mathcal{N}$, the contratensor product functor $\mathcal{N}\odot_{\mathfrak{R}}\nobreak{-}\,\colon\allowbreak\mathfrak{R}{\operatorname{\mathsf{--Contra}}}\longrightarrow\mathsf{Ab}$ restricted to the full subcategory $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}\subset\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ is exact on $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$. In particular, for any open right ideal $\mathfrak{I}\subset\mathfrak{R}$, the reduction functor $\mathfrak{F}\longmapsto\mathfrak{F}/(\mathfrak{I}\rightthreetimes\mathfrak{I})$ is exact on $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$.
(b) The directed colimit functors are exact in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ (but not in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$).
Proof.
Part (a) is [34, Lemma 8.4(b)]. Part (b) is [34, Lemma 8.8]. ∎
Lemma 4.4.
Let $\mathfrak{R}$ be a complete, separated topological ring, and let $\mathfrak{R}\supset\mathfrak{I}_{1}\supset\mathfrak{I}_{2}\supset\mathfrak{I}_{3}\supset\dotsb$ be a countable base of neighborhoods of zero in $\mathfrak{R}$ consisting of open two-sided ideals $\mathfrak{I}_{n}$,  $n\geq 1$. Denote by $R_{n}=\mathfrak{R}/\mathfrak{I}_{n}$ the corresponding discrete quotient rings. Then the rule assigning to a flat left $\mathfrak{R}$-contramodule $\mathfrak{F}$ the collection of $R_{n}$-modules $F_{n}=\mathfrak{F}/(\mathfrak{I}_{n}\rightthreetimes\mathfrak{F})$ is an equivalence between the category of flat left $\mathfrak{R}$-contramodules $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ and the category of the following sets of data:
  1. (1)
    for every $n\geq 1$, a flat left $R_{n}$-module $F_{n}$ is given;
  2. (2)
    for every $n\geq 1$, an isomorphism of left $R_{n}$-modules $F_{n}\simeq R_{n}\otimes_{R_{n+1}}F_{n+1}$ is given.
Proof.
This is [34, Lemma 8.6(b)]. ∎
Proposition 4.5.
Let $\kappa$ be a regular cardinal, and let $\mathfrak{R}$ be a complete, separated topological ring with a base of neighborhoods of zero of cardinality smaller than $\kappa$, consisting of open right ideals. Then the abelian category of left $\mathfrak{R}$-contramodules $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ is locally $\kappa$-presentable. The $\kappa$-presentable objects of $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ are precisely all the cokernels of morphisms of free $\mathfrak{R}$-contramodules spanned by sets of cardinality smaller than $\kappa$.
Proof.
This is [34, Proposition 9.1(a)]. ∎
Assuming that the topological ring $\mathfrak{R}$ has a countable base of neighborhoods of zero consisting of open right ideals, we will say that a left $\mathfrak{R}$-contramodule $\mathfrak{P}$ is countably generated if $\mathfrak{P}$ is a quotient contramodule of a free contramodule $\mathfrak{R}[[X]]$ spanned by a countable set $X$. This is equivalent to $\mathfrak{P}$ being an $\aleph_{1}$-generated object of $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ in the sense of [1, Definition 1.67].
Proposition 4.6.
Let $\mathfrak{R}$ be a complete, separated topological ring, and let $\mathfrak{R}\supset\mathfrak{I}_{1}\supset\mathfrak{I}_{2}\supset\mathfrak{I}_{3}\supset\dotsb$ be a countable base of neighborhoods of zero in $\mathfrak{R}$ consisting of open two-sided ideals $\mathfrak{I}_{n}$,  $n\geq 1$. Denote by $R_{n}=\mathfrak{R}/\mathfrak{I}_{n}$ the corresponding discrete quotient rings. Then the category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ of flat left $\mathfrak{R}$-contramodules is $\aleph_{1}$-accessible. Furthermore,
(a) a flat left $\mathfrak{R}$-contramodule is $\aleph_{1}$-presentable as an object of $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ if and only if it is $\aleph_{1}$-presentable as an object of $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$;
(b) a flat left $\mathfrak{R}$-contramodule $\mathfrak{F}$ is $\aleph_{1}$-presentable if and only if the $R_{n}$-module $F_{n}=\mathfrak{F}/(\mathfrak{I}_{n}\rightthreetimes\mathfrak{F})$ is countably presentable for every $n\geq 1$.
Proof.
This is [34, Lemmas 9.4 and 9.5, and Theorem 10.1]. ∎
Proposition 4.7.
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open two-sided ideals. Then any countably presentable flat $\mathfrak{R}$-contramodule has projective dimension $\leq\nobreak 1$ in the abelian category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$.
Proof.
This is [34, Corollary 9.7]. ∎
5. Acyclic Complexes of Flat Contramodules are Contraacyclic
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open right ideals. We will say that a complex of flat left $\mathfrak{R}$-contramodules $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is pure acyclic if it is acyclic in the exact category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$[34, Section 11]. Equivalently, a complex of flat $\mathfrak{R}$-contramodules $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is pure acyclic if and only if $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is acyclic in the abelian category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ and the $\mathfrak{R}$-contramodules of cocycles of $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ are flat. Let us denote the full subcategory of pure acyclic complexes of flat left $\mathfrak{R}$-contramodules by $\mathsf{Ac}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})\subset\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})\subset\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$.
The aim of this section is to prove the following theorem.
Theorem 5.1.
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open two-sided ideals. Then, for any complex of projective left $\mathfrak{R}$-contramodules $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ and any pure acyclic complex of flat left $\mathfrak{R}$-contramodules $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$, any morphism of complexes $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero. In other words, any pure acyclic complex of flat $\mathfrak{R}$-contramodules is contraacyclic (as a complex in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$, or equivalently, in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$).
We start with collecting some results from [40] and [34] on which the proof of Theorem 5.1 is based.
Proposition 5.2.
Let $\kappa$ be a regular cardinal and $\mathsf{B}$ be an abelian category with infinite coproducts and a $\kappa$-presentable projective generator $P$ (these conditions imply that the category $\mathsf{B}$ is locally $\kappa$-presentable). Denote by $\mathsf{S}$ the set of all (representatives of isomorphism classes of) bounded below complexes in $\mathsf{B}$ whose terms are coproducts of less then $\kappa$ copies of $P$. Then any complex of projective objects in $\mathsf{B}$ is a direct summand of a complex filtered by complexes from $\mathsf{S}$.
Proof.
Notice, first of all, that any projective object in $\mathsf{B}$ is a direct summand of a coproduct of copies of $P$; hence any complex of projective objects in $\mathsf{B}$ is a direct summand of a complex whose components are coproducts of copies of $P$. For the purposes of this paper, we are really interested in the case $\kappa=\aleph_{1}$. When $\kappa=\aleph_{0}$, the category $\mathsf{B}$ is the module category $\mathsf{B}\simeq R{\operatorname{\mathsf{--Mod}}}$, where $R$ is (the opposite ring to) the endomorphism ring of the small projective generator $P\in\mathsf{B}$. This case is covered by [44, Proposition 4.3] (cf. [26, Construction 4.3]). In the case of an uncountable regular cardinal $\kappa$, the desired assertion was established in [40, second proof of Proposition 7.2]. ∎
Proposition 5.3.
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open two-sided ideals. Then the category $\mathsf{Ac}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$ of pure acyclic complexes of flat left $\mathfrak{R}$-contramodules is $\aleph_{1}$-accessible. A pure acyclic complex of flat $\mathfrak{R}$-contramodules $\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is $\aleph_{1}$-presentable as an object of $\mathsf{Ac}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$ if and only if all the $\mathfrak{R}$-contramodules $\mathfrak{H}^{n}$,  $n\in\mathbb{Z}$, are countably presentable.
Proof.
This is [34, Corollary 11.4]. ∎
Proof of Theorem 5.1.
By Lemma 2.4, we have
$\operatorname{Hom}_{\mathsf{Hot}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})}(\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}})\simeq\operatorname{Ext}^{1}_{\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})}(\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}[-1]).$
Proposition 5.2 for $\mathsf{B}=\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$,  $\kappa=\aleph_{1}$, and $P=\mathfrak{R}=\mathfrak{R}[\{*\}]\in\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ tells us that $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is a direct summand of a complex filtered by complexes of free $\mathfrak{R}$-contramodules with (at most) countable sets of generators. In view of Lemma 2.2, the question reduces to the case when $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is a complex of free $\mathfrak{R}$-contramodules with countable sets of generators.
Then, by Propositions 1.2(a) and 4.5, the complex $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is a countably presentable object of the category $\mathsf{K}=\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$. On the other hand, by Proposition 5.3, the complex $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is an $\aleph_{1}$-directed colimit of pure acyclic complexes of countably presentable flat $\mathfrak{R}$-contramodules. Notice that the full subcategory of pure acyclic complexes of flat $\mathfrak{R}$-contamodules is closed under directed colimits in $\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$ by Lemma 4.3(b); so the directed colimits in this full subcategory agree with the ones in $\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$. Consequently, any morphism of complexes $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ factorizes through some pure acyclic complex of countably presentable flat $\mathfrak{R}$-contramodules $\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$.
It remains to recall that the homological dimension of the exact category of countably presentable flat $\mathfrak{R}$-contramodules does not exceed $1$ by Proposition 4.7. By Lemma 2.7, it follows that the complex $\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is absolutely acyclic in the exact category of countably presentable flat $\mathfrak{R}$-contramodules, and consequently also in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ and in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$. By Lemma 2.5(a), any morphism of complexes $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero. ∎
Remark 5.4.
Theorem 5.1 is a contramodule generalization of the similar assertion in Neeman’s paper [26, Theorem 8.6], which is the module version. Our proof of Theorem 5.1 is inspired by, but independent of the proof of the related assertion in [26].
Alternatively, it is possible to deduce our Theorem 5.1 from Neeman’s result. To this end, one argues as follows. Let $\mathfrak{R}\supset\mathfrak{I}_{1}\supset\mathfrak{I}_{2}\supset\mathfrak{I}_{3}\supset\dotsb$ be a countable base of neighborhoods of zero in $\mathfrak{R}$ consisting of open two-sided ideals $\mathfrak{I}_{n}$,  $n\geq 1$ (as in Lemma 4.4 and Proposition 4.6). Put $R_{n}=\mathfrak{R}/\mathfrak{I}_{n}$.
Then, by Lemma 4.4, for any flat left $\mathfrak{R}$-contramodules $\mathfrak{P}$ and $\mathfrak{F}$, there is a natural isomorphism of abelian groups
[…]
where the transition maps in the projective system are induced by the functors $R_{n}\otimes_{R_{n+1}}{-}\,\colon R_{n+1}{\operatorname{\mathsf{--Mod}}}\longrightarrow R_{n}{\operatorname{\mathsf{--Mod}}}$. When the $\mathfrak{R}$-contramodule $\mathfrak{P}$ is projective, the $R_{n}$-modules $\mathfrak{P}/\mathfrak{I}_{n}\rightthreetimes\mathfrak{P}$ are projective, too, and it follows that any $R_{n}$-module map $\mathfrak{P}/\mathfrak{I}_{n}\rightthreetimes\mathfrak{P}\longrightarrow\mathfrak{F}/\mathfrak{I}_{n}\rightthreetimes\mathfrak{F}$ can be lifted to an $R_{n+1}$-module map $\mathfrak{P}/\mathfrak{I}_{n+1}\rightthreetimes\mathfrak{P}\longrightarrow\mathfrak{F}/\mathfrak{I}_{n+1}\rightthreetimes\mathfrak{F}$. So the transition maps in the projective system are surjective.
Now, for any complex of projective left $\mathfrak{R}$-contramodules $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ and any complex of flat left $\mathfrak{R}$-contramodules $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$, we have a natural isomorphism of complexes of abelian groups
[…]
where the transition maps are surjective in every degree. Whenever the complex of flat $\mathfrak{R}$-contramodules $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is pure acyclic, so are the complexes of flat $R_{n}$-modules $\mathfrak{F}/\mathfrak{I}_{n}\rightthreetimes\mathfrak{F}$ (by Lemma 4.3(a)), so the complexes $\operatorname{Hom}_{R_{n}}^{\text{\smaller\smaller$\scriptstyle\bullet$}}(\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}/\mathfrak{I}_{n}\rightthreetimes\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\>\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}/\mathfrak{I}_{n}\rightthreetimes\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ are acyclic by [26, Theorem 8.6]. It remains to point out that the projective limit of a sequence (indexed by the positive integers) of termwise surjective maps of acyclic complexes of abelian groups is an acyclic complex again. Thus the complex $\operatorname{Hom}^{\mathfrak{R},{\text{\smaller\smaller$\scriptstyle\bullet$}}}(\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ is acyclic whenever the complex $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is pure acyclic.
6. Contraacyclic Complexes of Flat Contramodules are Acyclic
The aim of this section is to prove the following theorem, which provides the converse implication to Theorem 5.1.
Theorem 6.1.
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open two-sided ideals. Let $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be a complex of flat left $\mathfrak{R}$-contramodules. Assume that, for every complex of projective left $\mathfrak{R}$-contramodules $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$, all morphisms of complexes $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ are homotopic to zero. Then $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is a pure acyclic complex. In other words, any contraacyclic complex of flat $\mathfrak{R}$-contramodules is pure acyclic.
Once again, we need to collect some results from [38] and [34] on which the proof of Theorem 6.1 is based, and also prove some auxiliary lemmas. Let us start with the following definition.
A left $\mathfrak{R}$-contramodule $\mathfrak{C}$ is said to be cotorsion [38, Definition 7.3] if one has $\operatorname{Ext}^{\mathfrak{R},1}(\mathfrak{F},\mathfrak{C})=0$ for every flat left $\mathfrak{R}$-contramodule $\mathfrak{F}$. We denote the full subcategory of cotorsion left $\mathfrak{R}$-contramodules by $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}\subset\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$. The following lemma may help the reader to feel more comfortable.
Lemma 6.2.
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open right ideals. Then one has $\operatorname{Ext}^{\mathfrak{R},n}(\mathfrak{F},\mathfrak{C})=0$ for any flat left $\mathfrak{R}$-contramodule $\mathfrak{F}$, any cotorsion left $\mathfrak{R}$-contramodule $\mathfrak{C}$, and all $n\geq 1$.
Proof.
Compare Lemma 4.2 with Lemma 2.1. ∎
Proposition 6.3.
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open right ideals. Then every left $\mathfrak{R}$-contramodule $\mathfrak{M}$ can be included into a short exact sequence $0\longrightarrow\mathfrak{M}\longrightarrow\mathfrak{C}\longrightarrow\mathfrak{G}\longrightarrow 0$ in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ with a cotorsion $\mathfrak{R}$-contramodule $\mathfrak{C}$ and a flat $\mathfrak{R}$-contramodule $\mathfrak{G}$.
Proof.
This is a part of [38, Corollary 7.8] (see the terminology in [38, Section 3] or in Section 7 below). ∎
Lemma 6.4.
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open right ideals. Let $\mathfrak{G}$ be a left $\mathfrak{R}$-contramodule that is a quotient contramodule of a cotorsion $\mathfrak{R}$-contramodule. Then one has $\operatorname{Ext}^{\mathfrak{R},1}(\mathfrak{H},\mathfrak{G})=0$ for every flat left $\mathfrak{R}$-contramodule $\mathfrak{H}$ of projective dimension $1$.
Proof.
Let $0\longrightarrow\mathfrak{M}\longrightarrow\mathfrak{C}\longrightarrow\mathfrak{G}\longrightarrow 0$ be a short exact sequence in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ with a cotorsion $\mathfrak{R}$-contramodule $\mathfrak{C}$. Then we have a long exact sequence of abelian groups $\dotsb\longrightarrow\operatorname{Ext}^{\mathfrak{R},1}(\mathfrak{H},\mathfrak{C})\longrightarrow\operatorname{Ext}^{\mathfrak{R},1}(\mathfrak{H},\mathfrak{G})\longrightarrow\operatorname{Ext}^{\mathfrak{R},2}(\mathfrak{H},\mathfrak{M})\longrightarrow\dotsb$, implying the desired $\operatorname{Ext}^{\mathfrak{R},1}$ vanishing. ∎
Corollary 6.5.
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open right ideals. Let $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be a complex of left $\mathfrak{R}$-contramodules whose terms are quotient contramodules of cotorsion $\mathfrak{R}$-contramodules. Let $0\longrightarrow\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{Q}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow 0$ be a short exact sequence of complexes of left $\mathfrak{R}$-contramodules such that $\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is a complex of flat $\mathfrak{R}$-contramodules of projective dimension at most $1$. Denote by $\mathfrak{T}^{\text{\smaller\smaller$\scriptstyle\bullet$}}=\operatorname{Tot}(\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\to\mathfrak{Q}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\to\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ the related total complex. Then every morphism of complexes $\mathfrak{T}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero.
Proof.
In view of Lemma 6.4, we have a short exact sequence of complexes of abelian groups $0\longrightarrow\operatorname{Hom}^{\mathfrak{R},{\text{\smaller\smaller$\scriptstyle\bullet$}}}(\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow\operatorname{Hom}^{\mathfrak{R},{\text{\smaller\smaller$\scriptstyle\bullet$}}}(\mathfrak{Q}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow\operatorname{Hom}^{\mathfrak{R},{\text{\smaller\smaller$\scriptstyle\bullet$}}}(\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow 0$. So the dual argument to the proof of Lemma 2.5(a) applies, proving that the complex of abelian groups $\operatorname{Hom}^{\mathfrak{R},{\text{\smaller\smaller$\scriptstyle\bullet$}}}(\mathfrak{T}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ is acyclic. ∎
Corollary 6.6.
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open right ideals. Then every complex of flat left $\mathfrak{R}$-contramodules $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ can be included into a short exact sequence $0\longrightarrow\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow 0$ in $\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$ with a contractible complex of flat cotorsion $\mathfrak{R}$-contramodules $\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ and a complex of flat $\mathfrak{R}$-contramodules $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$.
Proof.
This is a corollary of Proposition 6.3. For every cohomological degree $n\in\mathbb{Z}$, choose a short exact sequence $0\longrightarrow\mathfrak{F}^{n}\longrightarrow\overline{\mathfrak{C}}^{n}\longrightarrow\overline{\mathfrak{G}}^{n}\longrightarrow 0$ in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ with a cotorsion $\mathfrak{R}$-contramodule ¯$\overline{\mathfrak{C}}^{n}$ and a flat $\mathfrak{R}$-contramodule $\overline{\mathfrak{G}}^{n}$. By Lemma 4.2, the $\mathfrak{R}$-contramodules ¯$\overline{\mathfrak{C}}^{n}$ are flat as extensions of flat $\mathfrak{R}$-contramodules $\mathfrak{F}^{n}$ and $\overline{\mathfrak{G}}^{n}$.
Let $\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be the direct sum of contractible two-term complexes $0\longrightarrow\overline{\mathfrak{C}}^{n}\overset{{\mathrm{id}}}{\longrightarrow}\overline{\mathfrak{C}}^{n}\longrightarrow 0$ with the cohomological grading defined by the rule that $\mathfrak{C}^{n}=\overline{\mathfrak{C}}^{n}\oplus\overline{\mathfrak{C}}^{n+1}$. Let $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be the morphism of complexes whose components are the monomorphisms $\mathfrak{F}^{n}\longrightarrow\overline{\mathfrak{C}}^{n}$ and the compositions $\mathfrak{F}^{n}\longrightarrow\mathfrak{F}^{n+1}\longrightarrow\overline{\mathfrak{C}}^{n+1}$. Then $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is a (termwise) monomorphism of complexes of $\mathfrak{R}$-contramodules. Let $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be the cokernel complex. Then there are short exact sequences of $\mathfrak{R}$-contramodules $0\longrightarrow\overline{\mathfrak{C}}^{n+1}\longrightarrow\mathfrak{G}^{n}\longrightarrow\overline{\mathfrak{G}}^{n}\longrightarrow 0$. Once again, Lemma 4.2 tells us that the $\mathfrak{R}$-contramodules $\mathfrak{G}^{n}$ are flat as extensions of flat $\mathfrak{R}$-contramodules. ∎
Proposition 6.7.
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open two-sided ideals. Then the category $\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$ of complexes of flat left $\mathfrak{R}$-contramodules is $\aleph_{1}$-accessible. A complex of flat $\mathfrak{R}$-contramodules $\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is $\aleph_{1}$-presentable as an object of $\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$ if and only if all the $\mathfrak{R}$-contramodules $\mathfrak{H}^{n}$,  $n\in\mathbb{Z}$, are countably presentable.
Proof.
This is [34, Proposition 10.2]. The assertion follows from Propositions 4.6(a) and 1.2(b) (with $\kappa=\aleph_{1}$ and $\lambda=\aleph_{0}$). ∎
Proof of Theorem 6.1.
Taking $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}=\mathfrak{R}[[\{*\}]]$ to be the one-term complex corresponding to the free $\mathfrak{R}$-contramodule with one generator $\mathfrak{R}[[\{*\}]]=\mathfrak{R}$, one can immediately see that any complex $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ satisfying the assumption of the theorem is acyclic in the abelian category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$. Our task is to prove that the complex $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is pure acyclic, i. e., acyclic in the exact category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$.
Let $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be a complex of flat left $\mathfrak{R}$-contramodules. Using Corollary 6.6, we can find a short exact sequence of complexes of $\mathfrak{R}$-contramodules $0\longrightarrow\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow 0$, where $\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is a contractible complex of flat cotorsion $\mathfrak{R}$-contramodules and $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is a complex of flat $\mathfrak{R}$-contramodules. Then the complex $\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$, being a contractible complex of flat contramodules, is both pure acyclic and contraacyclic.
Since the class of all acyclic complexes in an exact category is closed under kernels of termwise admissible epimorphisms and cokernels of termwise admissible monomorphisms, the complex $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is pure acyclic if and only if the complex $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is pure acyclic. On the other hand, in view of Lemma 2.5(a), the complex $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is contraacyclic if and only if the complex $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is contraacyclic. So we can consider $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ instead of $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ and assume that $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is contraacyclic. Our aim is to prove that $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is pure acyclic. What we have achieved here is that we know additionally that all the terms of the complex $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ are quotient contramodules of cotorsion contramodules.
[…]
It is easy to represent $\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ as a quotient complex of a complex of (countably generated) projective $\mathfrak{R}$-contramodules $\mathfrak{Q}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$. So we have a short exact sequence of complexes of $\mathfrak{R}$-contramodules $0\longrightarrow\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{Q}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow 0$. By Proposition 4.7,  $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is also a complex of projective $\mathfrak{R}$-contramodules (in fact, of countably generated projective $\mathfrak{R}$-contramodules, by [34, Lemma 11.1]). Put $\mathfrak{T}^{\text{\smaller\smaller$\scriptstyle\bullet$}}=\operatorname{Tot}(\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\to\mathfrak{Q}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\to\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$. Then the complex of abelian groups $\operatorname{Hom}^{\mathfrak{R},{\text{\smaller\smaller$\scriptstyle\bullet$}}}(\mathfrak{T}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ is acyclic by Corollary 6.5. The complexes of abelian groups $\operatorname{Hom}^{\mathfrak{R},{\text{\smaller\smaller$\scriptstyle\bullet$}}}(\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ and $\operatorname{Hom}^{\mathfrak{R},{\text{\smaller\smaller$\scriptstyle\bullet$}}}(\mathfrak{Q}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ are acyclic, since the complex $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is contraacyclic in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ by assumption. Thus the complex of abelian groups $\operatorname{Hom}^{\mathfrak{R},{\text{\smaller\smaller$\scriptstyle\bullet$}}}(\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ is acyclic, too.
We have shown that every morphism of complexes $f\colon\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero. Thus the morphism $f$ factorizes through the cone of the identity endomorphism of the complex $\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$. The latter cone is a contractible (hence pure acyclic) complex of countably presentable flat left $\mathfrak{R}$-contramodules. So it belongs to $\mathsf{T}$. ∎
7. Cotorsion Pairs
This section continues the discussion of category-theoretic background that was started in Section 2. Given an exact category $\mathsf{K}$, a pair of classes of objects $(\mathsf{F},\mathsf{C})$ in $\mathsf{K}$ is said to be a cotorsion pair if $\mathsf{C}=\mathsf{F}^{\perp_{1}}$ and $\mathsf{F}={}^{\perp_{1}}\mathsf{C}$.
For any class of objects $\mathsf{S}\subset\mathsf{K}$, the pair of classes $\mathsf{C}=\mathsf{S}^{\perp_{1}}$ and $\mathsf{F}={}^{\perp_{1}}\mathsf{C}$ is a cotorsion pair in $\mathsf{K}$. The cotorsion pair $(\mathsf{F},\mathsf{C})$ is said to be generated by the class of objects $\mathsf{S}$.
Dually to the definition in Section 2, a class of objects $\mathsf{C}\subset\mathsf{K}$ is said to be cogenerating if for every object $K\in\mathsf{K}$ there exists an object $C\in\mathsf{C}$ together with an admissible monomorphism $K\longrightarrow C$ in $\mathsf{K}$.
Let $(\mathsf{F},\mathsf{C})$ be a cotorsion pair in $\mathsf{K}$ such that the class $\mathsf{F}$ is generating and the class $\mathsf{C}$ is cogenerating in $\mathsf{K}$. Then the cotorsion pair $(\mathsf{F},\mathsf{C})$ is said to be hereditary if any one of the following equivalent conditions holds:
  1. (1)
    the class $\mathsf{F}$ is closed under kernels of admissible epimorphisms in $\mathsf{K}$;
  2. (2)
    the class $\mathsf{C}$ is closed under cokernels of admissible monomorphisms in $\mathsf{K}$;
  3. (3)
    $\operatorname{Ext}^{2}_{\mathsf{K}}(F,C)=0$ for all $F\in\mathsf{F}$ and $C\in\mathsf{C}$;
  4. (4)
    $\operatorname{Ext}^{n}_{\mathsf{K}}(F,C)=0$ for all $F\in\mathsf{F}$,  $C\in\mathsf{C}$, and $n\geq 1$.
The nontrivial implications are (1) $\Longrightarrow$ (4) and (2) $\Longrightarrow$ (4); they hold by Lemma 2.1 and its dual version.
A cotorsion pair $(\mathsf{F},\mathsf{C})$ is $\mathsf{K}$ is said to be complete if for every object $K\in\mathsf{K}$ there exist (admissible) short exact sequences
[…]
in $\mathsf{K}$ with objects $F$, $F^{\prime}\in\mathsf{F}$ and $C$, $C^{\prime}\in\mathsf{C}$. The short exact sequence (1) is said to be a special precover sequence. The short exact sequence (2) is said to be a special preenvelope sequence.
The following result is known classically as the Eklof–Trlifaj theorem [12, Theorems 2 and 10]. Here, given a class of objects $\mathsf{F}\subset\mathsf{K}$, we denote by $\mathsf{F}^{\oplus}\subset\mathsf{K}$ the class of all direct summands of the objects from $\mathsf{F}$. The notation $\mathsf{Fil}(\mathsf{S})\subset\mathsf{K}$ was introduced in Section 2.
Theorem 7.1.
Let $\mathsf{B}$ be a locally presentable abelian category and $\mathsf{S}\subset\mathsf{B}$ be a set of objects. Consider the cotorsion pair $(\mathsf{F},\mathsf{C})$ in $\mathsf{B}$ generated by $\mathsf{S}$. In this context:
(a) if the class $\mathsf{F}$ is generating in $\mathsf{B}$ and the class $\mathsf{C}$ is cogenerating in $\mathsf{B}$, then the cotorsion pair $(\mathsf{F},\mathsf{C})$ is complete;
(b) if the class $\mathsf{Fil}(\mathsf{S})$ is generating in $\mathsf{B}$, then $\mathsf{F}=\mathsf{Fil}(\mathsf{S})^{\oplus}$.
Proof.
Part (a) is [38, Corollary 3.6]. Part (b) is [38, Theorem 4.8(a,d)]. See also [40, Theorems 3.3 and 3.4]. ∎
The following result is closely related to Theorem 2.8.
Theorem 7.2.
Let $\mathsf{B}$ be a locally presentable abelian category with enough projective objects. Then the pair of classes of complexes of projective objects $\mathsf{F}=\mathsf{Com}(\mathsf{B}_{\mathsf{proj}})$ and contraacyclic complexes $\mathsf{C}=\mathsf{Ac}^{\mathsf{bctr}}(\mathsf{B})$ is a hereditary complete cotorsion pair $(\mathsf{F},\mathsf{C})$ in the abelian category of complexes $\mathsf{Com}(\mathsf{B})$.
Proof.
This is [40, Theorem 7.3]. For a generalization, see [41, Theorem 6.16]. ∎
Let $(\mathsf{F},\mathsf{C})$ be a complete cotorsion pair in $\mathsf{K}$ and $\mathsf{E}\subset\mathsf{K}$ be a full subcategory closed under extensions. We will say that the cotorsion pair $(\mathsf{F},\mathsf{C})$ restricts to a complete cotorsion pair in $\mathsf{E}$ if the pair of classes $\mathsf{E}\cap\mathsf{F}$ and $\mathsf{E}\cap\mathsf{C}$ is a complete cotorsion pair in (the inherited exact structure on) $\mathsf{E}$.
Lemma 7.3.
Let $\mathsf{K}$ be an exact category and $\mathsf{E}\subset\mathsf{K}$ be a full subcategory closed under extensions. Let $(\mathsf{F},\mathsf{C})$ be a complete cotorsion pair in $\mathsf{K}$. In this setting:
(a) if $\mathsf{E}$ is closed under kernels of admissible epimorphisms in $\mathsf{K}$ and $\mathsf{F}\subset\mathsf{E}$, then the cotorsion pair $(\mathsf{F},\mathsf{C})$ in $\mathsf{K}$ restricts to a complete cotorsion pair $(\mathsf{F},\,\mathsf{E}\cap\mathsf{C})$ in $\mathsf{E}$;
(b) if $\mathsf{E}$ is closed under cokernels of admissible monomorphisms in $\mathsf{K}$ and $\mathsf{C}\subset\mathsf{E}$, then the cotorsion pair $(\mathsf{F},\mathsf{C})$ in $\mathsf{K}$ restricts to a complete cotorsion pair $(\mathsf{E}\cap\mathsf{F},\,\mathsf{C})$ in $\mathsf{E}$.
Furthermore, in both cases the restricted cotorsion pair $(\mathsf{E}\cap\mathsf{F},\,\mathsf{E}\cap\mathsf{C})$ in $\mathsf{E}$ is hereditary whenever the cotorsion pair $(\mathsf{F},\mathsf{C})$ in $\mathsf{K}$ is hereditary.
Proof.
This is [33, Lemmas 1.5 and 1.6]. ∎
An exact category $\mathsf{E}$ is said to have exact directed colimits if all directed colimits exist in $\mathsf{E}$ and preserve (admissible) short exact sequences. The following important proposition goes back to [45, proofs of Lemma 5.2 and Proposition 5.3], [18, Propositions 3.1 and 3.2], and [2, proofs of Lemma 4.6 and Theorem 4.7].
Proposition 7.4.
Let $\mathsf{E}$ be an exact category with exact directed colimits and $\mathsf{A}\subset\mathsf{E}$ be a full subcategory closed under extensions and cokernels of admissible monomorphisms in $\mathsf{E}$. Then $\mathsf{A}$ is closed under transfinitely iterated extensions if and only if $\mathsf{A}$ is closed under directed colimits in $\mathsf{E}$.
Proof.
This is [42, Proposition 8.1]. ∎
In the following lemma, the class of objects $\mathsf{F}={}^{\perp_{1}}\mathsf{C}\subset\mathsf{K}$ is viewed as a full subcategory of $\mathsf{K}$ endowed with the exact structure inherited from $\mathsf{K}$.
Lemma 7.5.
Let $\mathsf{K}$ be an exact category and $\mathsf{C}\subset\mathsf{K}$ be a full subcategory. Put $\mathsf{F}={}^{\perp_{1}}\mathsf{C}\subset\mathsf{K}$. Let $C^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be a complex in $\mathsf{C}$. In this context:
(a) For any (termwise admissible) short exact sequence $0\longrightarrow F^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow G^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow H^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow 0$ of complexes in $\mathsf{F}$, the short sequence of complexes of abelian groups $0\longrightarrow\operatorname{Hom}_{\mathsf{K}}^{\text{\smaller\smaller$\scriptstyle\bullet$}}(H^{\text{\smaller\smaller$\scriptstyle\bullet$}},C^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow\operatorname{Hom}_{\mathsf{K}}^{\text{\smaller\smaller$\scriptstyle\bullet$}}(G^{\text{\smaller\smaller$\scriptstyle\bullet$}},C^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow\operatorname{Hom}_{\mathsf{K}}^{\text{\smaller\smaller$\scriptstyle\bullet$}}(F^{\text{\smaller\smaller$\scriptstyle\bullet$}},C^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow 0$ is exact.
(b) For any absolutely acyclic complex $A^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ in $\mathsf{F}$, any morphism of complexes $A^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow C^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero.
Proof.
This is a generalization of the dual version of Lemma 2.5(a). Part (a) is obvious; and to prove part (b), one can observe that acyclicity of the complex of abelian groups $\operatorname{Hom}_{\mathsf{K}}^{\text{\smaller\smaller$\scriptstyle\bullet$}}(A^{\text{\smaller\smaller$\scriptstyle\bullet$}},C^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ follows from part (a). ∎
In the next lemma, both the classes $\mathsf{F}$ and $\mathsf{C}\subset\mathsf{K}$ are viewed as full subcategories endowed with the exact structures inherited from $\mathsf{K}$.
Lemma 7.6.
Let $\mathsf{K}$ be an exact category and $(\mathsf{F},\mathsf{C})$ be a cotorsion pair in $\mathsf{K}$. Let $F^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be an acyclic complex in the exact category $\mathsf{F}$ and $C^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be an acyclic complex in the exact category $\mathsf{C}$. Then, viewing both $F^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ and $C^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ as complexes in $\mathsf{K}$, any morphism of complexes $F^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow C^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero.
Proof.
This is [15, Lemma 3.9] or [29, Lemma B.1.8]. ∎
8. Cotorsion Pair with Complexes of Cotorsion Contramodules
The aim of this section is to prove a contramodule version of the theorem of Bazzoni, Cortés-Izurdiaga, and Estrada [2, Theorem 5.3]. The argument in [2] is based on cotorsion periodicity together with [15, Corollary 4.10]. The proof of the latter uses considerations of deconstructibility in the category of complexes.
Such deconstructibility results have been established in full generality for Grothendieck abelian categories [44, Section 4], which does not cover our context. (Indeed, the directed colimits in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ are not exact, while $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ is not an abelian category.) Our proof in this section does not use that much knowledge about deconstructibility, but is based on accessibility considerations instead.
We start with the following proposition going back to [7, Lemma 1 and Proposition 2]. The definition of a deconstructible class was given in Section 2.
Proposition 8.1.
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open right ideals. Then the class of all flat left $\mathfrak{R}$-contramodules $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ is deconstructible in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$.
Proof.
This is [38, Corollary 7.6]; see also [37, Corollary 13.9] for a generalization. ∎
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open right ideals. The notation $\mathsf{Ac}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$ was introduced in Section 5, and the notation $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}$ in Section 6.
By [38, Corollary 7.8], the pair of classes $\mathsf{F}=\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ and $\mathsf{C}=\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}$ is a complete cotorsion pair $(\mathsf{F},\mathsf{C})$ in the abelian category $\mathsf{B}=\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ (cf. Proposition 6.3). By Lemma 4.2 or Lemma 6.2, this cotorsion pair is also hereditary. In the case of modules over an associative ring $R$, these results go back to [12, Theorems 2 and 10] and [7, Proposition 2 and Theorem 3].
The following Theorem 8.3 is our contramodule generalization of [2, Theorem 5.3] (see also [9, Theorem 3.3] or [29, Lemma 5.1.1(b)] for a quasi-coherent sheaf version). In order to prove it, we need more restrictive assumptions.
Remark 8.2.
Before formulating the theorem, let us say a few words about relevant references. In the paper of Bazzoni, Cortés-Izurdiaga, and Estrada [2], the assertion of [2, Theorem 5.3] is deduced from the cotorsion periodicity theorem [2, Theorem 5.1(2)] and a result of Gillespie [15, Corollary 4.10]. In the context of quasi-coherent sheaves, the assertion of [29, Lemma 5.1.1(b)] is likewise deduced from the cotorsion periodicity theorem for quasi-coherent sheaves. Both of these results for quasi-coherent sheaves were first stated by Christensen, Estrada, and Thompson [9, Theorem 3.3], but the proof contained a gap. In fact, the argument in [9, first paragraph of the proof of Lemma 3.2] is erroneous. Another proof of the cotorsion periodicity for quasi-coherent sheaves appeared in the paper [42, Theorem 10.2 and Corollary 10.4] by the present author and Št’ovíček, while the error in [9] was subsequently corrected by Estrada, Gillespie, and Odabaşi in [14, Theorem 6.3(2) and Remark 6.7]. See [42, Remark 10.3] for a discussion.
Theorem 8.3.
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open two-sided ideals. Then the pair of classes of pure acyclic complexes of flat contramodules $\mathsf{F}=\mathsf{Ac}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$ and arbitrary complexes of cotorsion contramodules $\mathsf{C}=\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})$ is a hereditary complete cotorsion pair $(\mathsf{F},\mathsf{C})$ in the abelian category of complexes of left $\mathfrak{R}$-contramodules $\mathsf{B}=\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$.
Proof.
By Proposition 8.1, there exists a set of flat left $\mathfrak{R}$-contramodules $\mathsf{S}_{0}\subset\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ such that $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}=\mathsf{Fil}(\mathsf{S}_{0})$. Denote by $\mathsf{S}\subset\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$ the set of all contractible two-term complexes of (flat) left $\mathfrak{R}$-contramodules
[…]
(situated in arbitrary cohomological degrees) with $\mathfrak{S}\in\mathsf{S}_{0}$. Then all contractible two-term complexes of flat left $\mathfrak{R}$-contamodules, and in particular, all contractible two-term complexes of projective left $\mathfrak{R}$-contramodules belong to $\mathsf{Fil}(\mathsf{S})$. By [38, Lemma 4.6(d)], we have $\mathsf{Fil}(\mathsf{Fil}(\mathsf{S}))=\mathsf{Fil}(\mathsf{S})$; hence the class $\mathsf{Fil}(\mathsf{S})$ is closed under coproducts in $\mathsf{B}$. So all contractible complexes of projective left $\mathfrak{R}$-contramodules belong to $\mathsf{Fil}(\mathsf{S})$. It follows easily that the class $\mathsf{Fil}(\mathsf{S})$ is generating in $\mathsf{B}$.
Let $(\mathsf{F}^{\prime},\mathsf{C}^{\prime})$ denote the cotorsion pair generated by $\mathsf{S}$ in $\mathsf{B}$. By Lemmas 2.2 and 2.3 (for $i=1$), we have $\mathsf{C}^{\prime}\subset\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})$. Conversely, by Lemma 2.4, $\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})\subset\mathsf{C}^{\prime}$. We have shown that $\mathsf{C}^{\prime}=\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})=\mathsf{C}$.
The construction from the proof of Corollary 6.6 (based on Proposition 6.3) shows that any complex of $\mathfrak{R}$-contramodules can be embedded as a subcomplex into a complex of cotorsion $\mathfrak{R}$-contramodules. So the class $\mathsf{C}^{\prime}$ is cogenerating in $\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$. Therefore, both parts of Theorem 7.1 are applicable, and we can conclude that $(\mathsf{F}^{\prime},\mathsf{C}^{\prime})$ is a complete cotorsion pair in $\mathsf{B}$ with $\mathsf{F}^{\prime}=\mathsf{Fil}(\mathsf{S})$.
Since the class of flat $\mathfrak{R}$-contramodules $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ is closed under extensions and directed colimits in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ (see Section 3 and Lemma 4.2), it is also closed under transfinitely iterated extensions. Therefore, we have $\mathsf{F}^{\prime}\subset\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$. Since the class of acyclic complexes in any exact category is closed under extensions, and the directed colimit functors are exact in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ by Lemma 4.3(b), it follows that, moreover, $\mathsf{F}^{\prime}\subset\mathsf{Ac}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$.
Let us show that the full subcategory $\mathsf{F}^{\prime}$ is closed under directed colimits in $\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$. Indeed, in order to prove that $\mathsf{F}^{\prime}$ is closed under cokernels of admissible monomorphisms in $\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$, consider a short exact sequence $0\longrightarrow\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow 0$ in $\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$. Then, for any complex $\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\in\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})$, we have a short exact sequence of complexes of abelian groups $0\longrightarrow\operatorname{Hom}^{\mathfrak{R},{\text{\smaller\smaller$\scriptstyle\bullet$}}}(\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow\operatorname{Hom}^{\mathfrak{R},{\text{\smaller\smaller$\scriptstyle\bullet$}}}(\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow\operatorname{Hom}^{\mathfrak{R},{\text{\smaller\smaller$\scriptstyle\bullet$}}}(\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow 0$ by Lemma 7.5(a). Therefore, if the complexes $\operatorname{Hom}^{\mathfrak{R},{\text{\smaller\smaller$\scriptstyle\bullet$}}}(\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ and $\operatorname{Hom}^{\mathfrak{R},{\text{\smaller\smaller$\scriptstyle\bullet$}}}(\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ are acyclic, then so is the complex $\operatorname{Hom}^{\mathfrak{R},{\text{\smaller\smaller$\scriptstyle\bullet$}}}(\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$. In view of Lemma 2.4, this means that one has $\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\in\mathsf{F}^{\prime}={}^{\perp_{1}}\mathsf{C}$ whenever $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\in\mathsf{F}^{\prime}$ and $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\in\mathsf{F}^{\prime}$. On the other hand, the full subcategory $\mathsf{F}^{\prime}$ is closed under transfinitely iterated extensions in $\mathsf{B}=\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$, hence also in $\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$, by Lemma 2.2. Applying Proposition 7.4, we come to the desired conclusion that $\mathsf{F}^{\prime}$ is closed under directed colimits in $\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$.
By Proposition 5.3, all the complexes belonging to $\mathsf{Ac}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$ are directed colimits of pure acyclic complexes of countably presentable flat $\mathfrak{R}$-contramodules. By Proposition 4.7 and Lemma 2.7, all pure acyclic complexes of countably presentable flat $\mathfrak{R}$-contramodules are absolutely acyclic as complexes in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$. By Lemmas 7.5(b) and 2.4, all absolutely acyclic complexes in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ belong to $\mathsf{F}^{\prime}={}^{\perp_{1}}\mathsf{C}$. Thus all pure acyclic complexes of countably presentable flat $\mathfrak{R}$-contramodules belong to $\mathsf{F}^{\prime}$. Using the result of the previous paragraph, we can conclude that $\mathsf{Ac}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})\subset\mathsf{F}^{\prime}$. We have shown that $\mathsf{F}^{\prime}=\mathsf{Ac}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})=\mathsf{F}$.
So we have constructed the desired complete cotorsion pair $(\mathsf{F},\mathsf{C})$ in $\mathsf{B}$. It remains to observe that the cotorsion pair $(\mathsf{F},\mathsf{C})$ is hereditary, because the class $\mathsf{C}$ is closed under cokernels of admissible monomorphisms in $\mathsf{B}$. This follows from the fact that the class $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}$ is closed under cokernels of admissible monomorphisms in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ (see the remarks preceding this theorem). ∎
Corollary 8.4.
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open two-sided ideals. Let $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be a complex of flat left $\mathfrak{R}$-contramodules. Then $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is a pure acyclic complex if and only if, for every complex of cotorsion left $\mathfrak{R}$-contramodules $\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$, all morphisms of complexes $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ are homotopic to zero.
Proof.
Follows from Theorem 8.3 in view of Lemma 2.4. ∎
9. Coderived Category of Flat Contramodules
We start with spelling out the definition dual to the one in Section 2. Let $\mathsf{E}$ be an exact category with enough injective objects. A complex $A^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ in $\mathsf{E}$ is said to be coacyclic (in the sense of Becker [4, Proposition 1.3.6(2)]) if, for every complex of injective objects $J^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ in $\mathsf{E}$, the complex of abelian groups $\operatorname{Hom}_{\mathsf{E}}^{\text{\smaller\smaller$\scriptstyle\bullet$}}(A^{\text{\smaller\smaller$\scriptstyle\bullet$}},J^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ is acyclic. The full subcategory of coacyclic complexes is denoted by $\mathsf{Ac}^{\mathsf{bco}}(\mathsf{E})\subset\mathsf{Hot}(\mathsf{E})$ or $\mathsf{Ac}^{\mathsf{bco}}(\mathsf{E})\subset\mathsf{Com}(\mathsf{E})$. The triangulated Verdier quotient category
$\mathsf{D}^{\mathsf{bco}}(\mathsf{E})=\mathsf{Hot}(\mathsf{E})/\mathsf{Ac}^{\mathsf{bco}}(\mathsf{E})$
is called the (Becker) coderived category of $\mathsf{E}$.
In this section, we are interested in the exact category of flat $\mathfrak{R}$-contramodules $\mathsf{E}=\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$. Consider the hereditary complete cotorsion pair $\mathsf{F}=\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ and $\mathsf{C}=\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}$ in the ambient abelian category $\mathsf{B}=\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ (as per the discussion in Section 8). Let us introduce the notation $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}}$ for the intersection of the two full subcategories
$\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}}=\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}\cap\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}\subset\mathfrak{R}{\operatorname{\mathsf{--Contra}}}.$
By Lemma 7.3(a), the cotorsion pair $(\mathsf{F},\mathsf{C})$ in $\mathsf{B}$ restricts to a hereditary complete cotorsion pair ($\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$, $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}})$ in the exact subcategory $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}=\mathsf{E}\subset\mathsf{B}$. In other words, this means that there are enough injective objects in the exact category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$, and these are precisely all the flat cotorsion $\mathfrak{R}$-contramodules.
Corollary 9.1.
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open two-sided ideals. Let $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be a complex of flat left $\mathfrak{R}$-contramodules. Then $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is a pure acyclic complex if and only if, for every complex of flat cotorsion left $\mathfrak{R}$-contramodules $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$, all morphisms of complexes $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ are homotopic to zero. In other words, a complex is coacyclic in the exact category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ if and only if it is acyclic in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$.
Proof.
Consider the abelian category $\mathsf{B}=\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$ and the hereditary complete cotorsion pair $\mathsf{F}=\mathsf{Ac}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$ and $\mathsf{C}=\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})$ in $\mathsf{B}$, as per Theorem 8.3. By Lemma 7.3(a), the cotorsion pair $(\mathsf{F},\mathsf{C})$ in $\mathsf{B}$ restricts to a hereditary complete cotorsion pair $(\mathsf{F},\,\mathsf{E}\cap\mathsf{C})$ in the exact subcategory $\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})=\mathsf{E}\subset\mathsf{B}$. Now $\mathsf{E}\cap\mathsf{C}=\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}})$ is the full subcategory of all complexes of flat cotorsion $\mathfrak{R}$-contramodules. So we have $\mathsf{F}={}^{\perp_{1}}(\mathsf{E}\cap\mathsf{C})$ in the exact category $\mathsf{E}$. In view of Lemma 2.4, the assertion of the corollary follows. ∎
Corollary 9.2.
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open two-sided ideals. Then the inclusion of additive/exact categories $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ induces a triangulated equivalence
$\mathsf{Hot}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}})\simeq\mathsf{D}^{\mathsf{bco}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})=\mathsf{D}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}).$
Proof.
Corollary 9.1 tells us that $\mathsf{D}^{\mathsf{bco}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})=\mathsf{D}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$. It follows immediately from the definitions that the functor $\mathsf{Hot}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}})\longrightarrow\mathsf{D}^{\mathsf{bco}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$ is fully faithful. In other to prove that it is a triangulated equivalence, it remains to construct for every complex of flat $\mathfrak{R}$-contramodules $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ a complex of flat cotorsion $\mathfrak{R}$-contramodules $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ together with a morphism of complexes $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ with a pure acyclic cone.
For this purpose, we use Theorem 8.3. As explained in the proof of Corollary 9.1, the theorem implies that the pair of classes $\mathsf{Ac}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$ and $\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}})$ is a hereditary complete cotorsion pair in $\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$. Consider a special preenvelope sequence (2) from Section 7 for the object $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\in\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$ with respect to this complete cotorsion pair. We get a short exact sequence of complexes of flat $\mathfrak{R}$-contramodules $0\longrightarrow\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow 0$ with $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\in\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}})$ and $\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\in\mathsf{Ac}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$. Since the total complex $\operatorname{Tot}(\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\to\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\to\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ is pure acyclic, it follows that the cone of the morphism of complexes $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ also belongs to $\mathsf{Ac}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$. ∎
10. Contraderived Category of Cotorsion Contramodules
The aim of this section is to prove a contramodule generalization of [36, Corollary 7.21]. The argument in [36, Section 7.7] for the existence of the relevant cotorsion pair uses considerations of deconstructibility in the category of complexes. As in Section 8 above, our proof in this section partly substitutes information about deconstructibility by an accessibility argument.
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open right ideals. We are interested in the exact category of cotorsion $\mathfrak{R}$-contramodules $\mathsf{E}=\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}$ (with the exact structure inherited from the abelian exact structure on $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$). Consider the hereditary complete cotorsion pair $\mathsf{F}=\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ and $\mathsf{C}=\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}$ in the abelian category $\mathsf{B}=\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$. By Lemma 7.3(b), the cotorsion pair $(\mathsf{F},\mathsf{C})$ in $\mathsf{B}$ restricts to a hereditary complete cotorsion pair ($\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}}$, $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}$) in the exact category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}=\mathsf{E}\subset\mathsf{B}$. In other words, this means that there are enough projective objects in the exact category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}$, and these are precisely all the flat cotorsion $\mathfrak{R}$-contramodules.
In the following theorem, which is our contramodule generalization of a construction of complete cotorsion pair from [36, proof of Theorem 7.19], we need more restrictive assumptions. (See also [29, Lemma 5.1.9(b)] for a quasi-coherent sheaf version, and [16, Proposition 3.2, Theorem 5.5, and Section 5.3] or [17, Lemma 4.9] for an abstract approach covering Grothendieck categories.)
Theorem 10.1.
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open two-sided ideals. Then the pair of classes of arbitrary complexes of flat contramodules $\mathsf{F}=\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$ and complexes of cotorsion contramodules contraacyclic as complexes in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$, i. e., $\mathsf{C}=\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})\cap\mathsf{Ac}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$, is a hereditary complete cotorsion pair $(\mathsf{F},\mathsf{C})$ in the abelian category of complexes of left $\mathfrak{R}$-contramodules $\mathsf{B}=\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$.
Proof.
The argument bears some similarity to the proof of Theorem 8.3. By Proposition 8.1, there exists a set of flat $\mathfrak{R}$-contramodules $\mathsf{S}_{0}\subset\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ such that $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}=\mathsf{Fil}(\mathsf{S}_{0})$. Denote by $\mathsf{S}_{1}\subset\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$ the set of all contractible two-term complexes of (flat) $\mathfrak{R}$-contramodules
[…]
(situated in arbitrary cohomological degrees) with $\mathfrak{S}\in\mathsf{S}_{0}$. Let $\mathsf{S}_{2}\subset\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$ be the set of (representatives of the isomorphism classes of) all complexes of countably presentable flat $\mathfrak{R}$-contramodules. By Proposition 6.7, all the complexes of flat $\mathfrak{R}$-contramodules are directed colimits of complexes from $\mathsf{S}_{2}$. Put $\mathsf{S}=\mathsf{S}_{1}\cup\mathsf{S}_{2}$. As explained in the proof of Theorem 8.3, the class $\mathsf{Fil}(\mathsf{S})$ (and even $\mathsf{Fil}(\mathsf{S}_{1})$) is generating in $\mathsf{B}$.
Let $(\mathsf{F}^{\prime},\mathsf{C}^{\prime})$ denote the cotorsion pair generated by $\mathsf{S}$ in $\mathsf{B}$. By Lemmas 2.2 and 2.3 (for $i=1$), we have $\mathsf{C}^{\prime}\subset\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})$. Conversely, by Lemma 2.4, all contractible complexes of cotorsion left $\mathfrak{R}$-contramodules belong to $\mathsf{C}^{\prime}$.
The construction from the proof of Corollary 6.6 (based on Proposition 6.3) shows that any complex of $\mathfrak{R}$-contramodules can be embedded as a subcomplex into a contractible complex of cotorsion $\mathfrak{R}$-contramodules. So the class $\mathsf{C}^{\prime}$ is cogenerating in $\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$. Therefore, both parts of Theorem 7.1 are applicable, and we have shown that $(\mathsf{F}^{\prime},\mathsf{C}^{\prime})$ is a complete cotorsion pair in $\mathsf{B}$ with $\mathsf{F}^{\prime}=\mathsf{Fil}(\mathsf{S})$. Similarly to the proof of Theorem 8.3, it follows that $\mathsf{F}^{\prime}\subset\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$.
Continuing to argue similarly to the proof of Theorem 8.3, one shows that the full subcategory $\mathsf{F}^{\prime}$ is closed under directed colimits in $\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$. Therefore, $\mathsf{F}^{\prime}=\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})=\mathsf{F}$ (since $\mathsf{S}_{2}\subset\mathsf{F}^{\prime}$). In particular, all the complexes of projective left $\mathfrak{R}$-contramodules belong to $\mathsf{F}^{\prime}$. By the definition and in view of Lemma 2.4, it follows that $\mathsf{C}^{\prime}\subset\mathsf{Ac}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$.
Conversely, let $\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be a complex of cotorsion $\mathfrak{R}$-contramodules that is contraacyclic as a complex in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$. Let $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be a complex of flat left $\mathfrak{R}$-contramodules. Let us show that $\operatorname{Ext}^{1}_{\mathsf{B}}(\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}})=0$. To this end, we will need to recall yet another complete cotorsion pair.
According to Theorem 7.2, the pair of classes ($\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{proj}})$, $\mathsf{Ac}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$) is a hereditary complete cotorsion pair in $\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$. By Lemma 7.3(a), this cotorsion pair restricts to a hereditary complete cotorsion pair in the exact subcategory $\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})\subset\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$ formed by the pair of classes $\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{proj}})$ and $\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})\cap\mathsf{Ac}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$. According to Theorems 5.1 and 6.1, we have $\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})\cap\mathsf{Ac}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})=\mathsf{Ac}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$. We have shown that the pair of classes ($\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{proj}})$, $\mathsf{Ac}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$) is a hereditary complete cotorsion pair in $\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$.
For our purposes, we need either a special precover sequence (1), or a special preenvelope sequence (2) for the complex $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ with respect to this cotorsion pair in $\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$. Both are equally suitable. For example, let $0\longrightarrow\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow 0$ be a special precover sequence; so $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\in\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{proj}})$ and $\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\in\mathsf{Ac}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$. Since $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\in\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$ and $\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\in\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})$, we have a short exact sequence of complexes of abelian groups $0\longrightarrow\operatorname{Hom}^{\mathfrak{R},{\text{\smaller\smaller$\scriptstyle\bullet$}}}(\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow\operatorname{Hom}^{\mathfrak{R},{\text{\smaller\smaller$\scriptstyle\bullet$}}}(\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow\operatorname{Hom}^{\mathfrak{R},{\text{\smaller\smaller$\scriptstyle\bullet$}}}(\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow 0$.
Now the complex $\operatorname{Hom}^{\mathfrak{R},{\text{\smaller\smaller$\scriptstyle\bullet$}}}(\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ is acyclic since $\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\in\mathsf{Ac}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$. Furthermore, the complex $\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is acyclic in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ by Lemma 2.6, hence it is also acyclic in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}$ by the cotorsion periodicity theorem for contramodules [34, Corollary 12.4]. As the complex $\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is acyclic in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$, it follows by virtue of Lemma 7.6 that the complex $\operatorname{Hom}^{\mathfrak{R},{\text{\smaller\smaller$\scriptstyle\bullet$}}}(\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ is acyclic. Therefore, the complex $\operatorname{Hom}^{\mathfrak{R},{\text{\smaller\smaller$\scriptstyle\bullet$}}}(\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ is acyclic, too. Applying Lemma 2.4, we conclude that $\operatorname{Ext}^{1}_{\mathsf{B}}(\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}})=0$. Thus $\mathsf{C}^{\prime}=\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})\cap\mathsf{Ac}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})=\mathsf{C}$.
We have obtained the desired complete cotorsion pair $(\mathsf{F},\mathsf{C})$ in $\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$, and it only remains to show that it is hereditary. Indeed, the class $\mathsf{F}$ is closed under kernels of admissible epimorphisms in $\mathsf{B}$ by Lemma 4.2. ∎
Corollary 10.2.
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open two-sided ideals. Let $\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be a complex of cotorsion left $\mathfrak{R}$-contramodules. Then the following two conditions are equivalent:
  1. (1)
    for every complex of projective left $\mathfrak{R}$-contramodules $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$, any morphism of complexes $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero;
  2. (2)
    for every complex of flat cotorsion left $\mathfrak{R}$-contramodules $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$, any morphism of complexes $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero.
In other words, $\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is contraacyclic as a complex in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ if and only if $\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is contraacyclic as a complex in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}$.
Proof.
Consider the abelian category $\mathsf{B}=\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$ and the hereditary complete cotorsion pair $\mathsf{F}=\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$ and $\mathsf{C}=\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})\cap\mathsf{Ac}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$ in $\mathsf{B}$, as per Theorem 10.1. By Lemma 7.3(b), the cotorsion pair $(\mathsf{F},\mathsf{C})$ in $\mathsf{B}$ restricts to a hereditary complete cotorsion pair $(\mathsf{E}\cap\mathsf{F},\,\mathsf{C})$ in the exact subcategory $\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})=\mathsf{E}\subset\mathsf{B}$. Now $\mathsf{E}\cap\mathsf{F}=\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}})$ is the full subcategory of all complexes of flat cotorsion $\mathfrak{R}$-contramodules. So we have $\mathsf{C}=(\mathsf{E}\cap\mathsf{F})^{\perp_{1}}$ in the exact category $\mathsf{E}$. In view of Lemma 2.4, the assertion of the corollary follows. ∎
Corollary 10.3.
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open two-sided ideals. Then the inclusion of additive/exact categories $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}$ induces a triangulated equivalence
$\mathsf{Hot}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}})\simeq\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}).$
Proof.
Corollary 10.2 tells us that $\mathsf{Ac}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})=\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})\cap\mathsf{Ac}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$. It follows immediately from the definitions that the functor $\mathsf{Hot}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}})\longrightarrow\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})$ is fully faithful. In order to prove that it is a triangulated equivalence, it remains to construct for every complex of cotorsion $\mathfrak{R}$-contramodules $\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ a complex of flat cotorsion $\mathfrak{R}$-contramodules $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ together with a morphism of complexes $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ with a contraacyclic cone.
For this purpose, we use Theorem 10.1. As explained in the proof of Corollary 10.2, the theorem implies that the pair of classes $\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}})$ and $\mathsf{Ac}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})$ is a hereditary complete cotorsion pair in $\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})$. Consider a special precover sequence (1) for the object $\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\in\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})$ with respect to this complete cotorsion pair. We get a short exact sequence of complexes of cotorsion $\mathfrak{R}$-contramodules $0\longrightarrow\mathfrak{B}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow 0$ with $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\in\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}})$ and $\mathfrak{B}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\in\mathsf{Ac}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})$. Since the total complex $\operatorname{Tot}(\mathfrak{B}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\to\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\to\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ is contraacyclic in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}$ by Lemma 2.5(a), it follows that the cone of the morphism of complexes $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ also belongs to $\mathsf{Ac}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})$. ∎
11. Five Constructions of the Contraderived Category
Now we can formulate the theorem promised in the introduction, summarizing our results generalizing Neeman’s theorem [26, Theorem 8.6] and the related implications of the theorem of Bazzoni, Cortés-Izurdiaga, and Estrada [2, Theorem 5.3].
Theorem 11.1.
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open two-sided ideals. Let $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be a complex of flat left $\mathfrak{R}$-contramodules. Then the following six conditions are equivalent:
  1. (ic)
    for every complex of projective left $\mathfrak{R}$-contamodules $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$, any morphism of complexes of $\mathfrak{R}$-contramodules $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero;
  2. (iic)
    $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is an acyclic complex of $\mathfrak{R}$-contramodules with flat $\mathfrak{R}$-contramodules of cocycles;
  3. (iiic)
    $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ can be obtained from contractible complexes of flat $\mathfrak{R}$-contramodules using extensions and directed colimits;
  4. (ivc)
    $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is an $\aleph_{1}$-directed colimit of total complexes of short exact sequences of complexes of countably presentable flat $\mathfrak{R}$-contramodules;
  5. (vc)
    for every complex of cotorsion left $\mathfrak{R}$-contramodules $\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$, any morphism of complexes of $\mathfrak{R}$-contramodules $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero;
  6. (vic)
    for every complex of flat cotorsion left $\mathfrak{R}$-contramodules $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$, any morphism of complexes of $\mathfrak{R}$-contramodules $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero.
Proof.
The implication ($\Longrightarrow$ (ic) is Theorem 5.1.
The implication ($\Longrightarrow$ (iic) is Theorem 6.1.
The equivalences (iic) $\Longleftrightarrow$ (iiic) $\Longleftrightarrow$ (ivc) are provided by [34, Theorem 13.2].
The equivalence (iic) $\Longleftrightarrow$ (vc) is Corollary 8.4.
The equivalence (iic) $\Longleftrightarrow$ (vic) is Corollary 9.1. ∎
Finally, we deduce a corollary about an equivalence of the “flat”, “projective”, “flat cotorsion”, “Verdier quotient by contraacyclic”, and “Verdier quotient of cotorsion by contraacyclic” constructions of the contraderived category $\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$. It should be compared with [43, Theorems 4.12, 4.13, and 5.7].
Corollary 11.2.
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open two-sided ideals. Then the inclusions of additive/exact/abelian categories $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{proj}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ induce equivalences of triangulated categories
$\mathsf{Hot}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{proj}})\simeq\mathsf{D}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})=\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})\simeq\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}).$
Furthermore, the inclusions of additive/exact/abelian categories $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ and $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ induce equivalences of triangulated categories
$\mathsf{Hot}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}})\simeq\mathsf{D}^{\mathsf{bco}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})=\mathsf{D}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$
and
$\mathsf{Hot}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}})\simeq\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})\simeq\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}).$
Proof.
The functor $\mathsf{Hot}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{proj}})\longrightarrow\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$ is a triangulated equivalence by Theorem 2.8. In fact, Theorem 2.8 provides more information than it explicitly says. For every complex of left $\mathfrak{R}$-contramodules $\mathfrak{M}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$, there exists a complex of projective left $\mathfrak{R}$-contramodules $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ together with a morphism of complexes $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{M}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ with a contraacyclic cone.
The assertion that the functor $\mathsf{D}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{proj}})\longrightarrow\mathsf{D}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$ is a triangulated equivalence is a corollary of Theorem 11.1 (ic) $\Leftrightarrow$ (iic). We apply Lemma 2.9. Consider a complex of flat left $\mathfrak{R}$-contramodules $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$. Then, as a particular case of the previous paragraph, there exists a complex of projective left $\mathfrak{R}$-contramodules $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ together with a morphism of complexes $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ with a contraacyclic cone. The cone, being a contraacyclic complex of flat $\mathfrak{R}$-contramodules, is consequently an acyclic complex in the exact category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$. So $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is an isomorphism in $\mathsf{D}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$. Furthermore, every complex of projective $\mathfrak{R}$-contramodules that is acyclic in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ is contractible, again by Theorem 11.1 (ic) $\Leftrightarrow$ (iic) or by [34, Proposition 12.1]. The desired conclusion follows.
One has $\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})=\mathsf{D}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})=\mathsf{D}^{\mathsf{bco}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$ by Theorem 11.1 (ic) $\Leftrightarrow$ (iic) $\Leftrightarrow$ (vic). The triangulated equivalence $\mathsf{Hot}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}})\simeq\mathsf{D}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$ is the result of Corollary 9.2. The triangulated equivalence $\mathsf{Hot}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}})\simeq\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})$ is the result of Corollary 10.3.
Finally, the triangulated functor $\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})\longrightarrow\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$ induced by the inclusion of exact/abelian categories $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ is well-defined by Corollary 10.2. This functor is a triangulated equivalence, since both the composition $\mathsf{Hot}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}})\longrightarrow\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})\longrightarrow\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$ and the functor $\mathsf{Hot}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}})\longrightarrow\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})$ are triangulated equivalences. ∎
Question 11.3.
It would be very interesting to extend the result of Theorem 11.1 (or some parts of it) to complete, separated topological rings $\mathfrak{R}$ with a countable base of neighborhoods of zero consisting of open right ideals. Is such a generalization true? See [34, Questions 9.8, 11.2, and 13.3] for further discussion.